### v1

# Evaluating Visual and Cultural Interpretation: The K-Viscuit Benchmark with Human-VLM Collaboration

Abstract
To create culturally inclusive vision-language models (VLMs), the foremost requirement is developing a test benchmark that can diagnose the models’ ability to respond to questions reflecting cultural elements. This paper addresses the necessity for such benchmarks, noting that existing research has relied on human annotators’ manual efforts, which impedes diversity and efficiency. We propose a semi-automated pipeline for constructing cultural VLM benchmarks to enhance diversity and efficiency. This pipeline leverages human-VLM collaboration, where VLMs generate questions based on guidelines, human-annotated examples, and image-wise relevant knowledge, which are then reviewed by native speakers for quality and cultural relevance. The effectiveness of our adaptable pipeline is demonstrated through a specific application: creating a dataset tailored to Korean culture, dubbed K-Viscuit. The resulting benchmark features two types of questions: Type 1 questions measure visual recognition abilities, while Type 2 assess fine-grained visual reasoning skills. This ensures a thorough diagnosis of VLM models across various aspects. Our evaluation using K-Viscuit revealed that open-source models notably lag behind proprietary models in understanding Korean culture, highlighting areas for improvement. We provided diverse analyses of VLM performance across different cultural aspects. Besides, we explored the potential of incorporating external knowledge retrieval to enhance the generation process, suggesting future directions for improving cultural interpretation ability of VLMs. Our dataset and code will be made publicly available.
Evaluating Visual and Cultural Interpretation:
The
K-Viscuit Benchmark with Human-VLM Collaboration
Yujin Baek11footnotemark: 11     ChaeHun Parkthanks:    Equal contribution1    Jaeseok Kim2     Yu-Jung Heo2    Du-Seong Chang2     Jaegul Choo1 1 KAIST AI     2 KT Corporation {yujinbaek,ddehun,jchoo}@kaist.ac.kr {jaeseok.kim,yj.heo,dschang}@kt.com
1 Introduction
The utility of vision-language models (VLMs) in bridging visual and textual data is well-recognized, offering advancements in applications such as image captioning and optical character recognition. However, most VLMs are predominantly trained on data from Western languages and cultures (Lin et al., 2014; Young et al., 2014; Antol et al., 2015), limiting their effectiveness in non-Western contexts (Liu et al., 2021; Yin et al., 2023, 2021; Romero et al., 2024). Cultural nuances significantly influence the interpretation of visual and textual information, impacting the accuracy and relevance of VLM outputs. Consequently, building VLMs that understand diverse cultural backgrounds is essential for enhancing inclusivity and ensuring fair performance across different demographics.
To develop VLMs that effectively understand the cultural contexts of specific countries, establishing a wide-ranging test benchmark is a critical first step. Such a benchmark is essential for diagnosing the model’s ability to accurately interpret and respond to culturally specific inputs. Notably, previous research has endeavored to create VLM test benchmarks for particular countries (Liu et al., 2021; Yin et al., 2021; Romero et al., 2024).
Figure 1: Pipeline Overview.
However, these approaches have primarily relied on manually crafted questions, which present two major limitations. First, their process is time-consuming and labor-intensive, increasing the cost of the research and making it difficult to extend and apply to the cultural contexts of other countries. Additionally, the diversity of questions can be limited due to cognitive fixation, where humans find it challenging to generate a wide range of varied questions after settling on an initial idea (Ramos, 2020). This lack of diversity can hinder the accurate diagnosis of the model’s ability to generalize across different types of culturally nuanced inputs.
To address these limitations, we propose a semi-automated pipeline for constructing cultural VLM benchmarks that enhance the efficiency and diversity of culture-relevant visual question and answer generation as illustrated in Fig. 1.
This pipeline incorporates human-VLM collaboration, where the VLM generates and recommends questions and answers leveraging carefully crafted guidelines, a small set of human-annotated examples, and image-wise relevant knowledge. These recommendations are then assessed and approved by native speakers, ensuring high quality and cultural relevance.
In this work, we applied the proposed pipeline to create a dataset for Korean culture, named K-Viscuit (Korean Visual and Cultural Interpretation Test). Nevertheless, our pipeline can be adapted and applied to other cultures analogously. The resulting benchmark dataset created using our proposed pipeline has two notable features. Firstly, it has the ability to measure nuanced cultural understanding through two types of assessments: Type 1, which measures visual recognition capability, and Type 2, which measures visual reasoning capability. Secondly, the multiple-choice questions in the dataset include four options, with distractors generated following carefully designed instructions to ensure high similarity among the options. This high similarity between options prevents models from using shortcuts to determine the correct answer, ensuring a more robust and comprehensive evaluation of cultural understanding in VLMs. These features underscore the effectiveness and thoroughness of the resulting benchmark dataset in addressing cultural nuances.
Using our benchmark data, we evaluated the performance of various open-source and proprietary VLMs. Our experimental results indicate that open-source models significantly lag behind proprietary models in the context of Korean culture. This assessment highlights key improvement areas and sets development goals for open-source models, with proprietary models as the reference. Our contributions are summarized as follows:
  • We propose a pipeline for building benchmarks to diagnose VLMs’ understanding of various cultures. It enables efficient creation of culturally diverse datasets through human-VLM cooperation.
  • Using our pipeline, we constructed the K-Viscuit Korean Culture VQA benchmark.
  • We utilized K-Viscuit to diagnose and analyze the performance of existing open-source and proprietary VLMs.
2 Related Work
Recent research has made significant strides in developing benchmarks to assess the ability of AI models, particularly VLMs and Large Language Models (LLMs), to understand and respond appropriately to various cultural contexts. These efforts are crucial given that many existing models are trained predominantly on Western-centric datasets (Young et al., 2014; Lin et al., 2014; Antol et al., 2015), which constrains their performance in non-Western settings. On the side of LLMs, Kim et al. (2024) introduced a benchmark dataset, CLIcK, specifically for assessing cultural and linguistic intelligence in Korean language models. This dataset includes QA pairs designed to test the cultural knowledge of Korean language models. Wibowo et al. (2023) created COPAL-ID, a dataset that captures Indonesian cultural nuances for text-based commonsense reasoning, highlighting the need for culturally aware benchmarks.
VLMs require consideration of both image and text inputs to effectively reflect cultural contexts, as cultural nuances can arise from either source. Liu et al. (2021) proposed the MaRVL dataset, a multilingual visually grounded reasoning dataset that includes five different languages and cultures to diagnose inherent language and cultural biases in pre-trained models. Their work highlights the need for models to seamlessly interpret and integrate visual and textual information across diverse cultures, paving the way for more inclusive and globally aware AI systems. Yin et al. (2021) built the GD-VCR dataset, which tests visual commonsense reasoning influenced by geographical and cultural contexts. Romero et al. (2024) presented the CVQA, a multilingual visual question-answering (VQA) benchmark that incorporates various languages and cultures, providing a comprehensive tool for evaluating VLMs in diverse cultural contexts.
Previous efforts to create cultural benchmarks for VLMs, however, have heavily relied on human annotators, which can limit both the diversity and efficiency of dataset creation (Liu et al., 2021; Yin et al., 2021; Ramaswamy et al., 2024; Romero et al., 2024). The efficiency and diversity of dataset creation can be significantly enhanced by leveraging AI models, particularly when combined with human collaboration. This approach not only improves the speed and scope of dataset generation but also maintains high-quality standards. Liu et al. (2022) demonstrated this in their study on building a natural language inference dataset, where they combined the generative capabilities of language models with human review. Analogous to how LLMs are heavily utilized in the creation of datasets for language-only tasks (Taori et al., 2023; Kim et al., 2023; Dubois et al., 2024; Kim et al., 2024), leveraging the inherent knowledge of AI models in building multimodal cultural benchmarks can yield diverse and culturally nuanced examples. In this work, by harnessing AI models, we have significantly enhanced the diversity and richness of the dataset, ensuring it better reflects the wide array of cultural contexts we aim to represent.
3 Data Construction Pipeline
We present our human-AI collaborated dataset construction pipeline to diagnose the understanding of VLMs in a target cultural field. In this work, we focus on Korean culture as a target cultural field. First, we provide an overview (section 3.1) and details of the pipeline. Then, we present the analysis of the resulting dataset, dubbed K-Viscuit (Korean Visual and Cultural Interpretation Test) (section 3.2).
3.1 Pipeline Overview
The pipeline is designed to comprise a multiple-choice visual question answering (VQA) task, where each evaluation sample consists of an image, a question, and four options with one correct choice. Native Korean speakers are involved in the dataset construction, and a powerful proprietary VLM is employed to mitigate unintended human biases, such as cognitive fixation (Ramos, 2020), and encourage an effective annotation process. The generated samples are designed to cover various aspects of a target culture derived from daily life and necessitate multimodal reasoning to interpret both the image and the question accurately. All text in the dataset is written in English to isolate the evaluation of multicultural comprehension from multilingual aspects. The dataset construction consists of four stages: 1) concept selection, 2) image selection, 3) question and options annotation, and 4) human verification. Fig. 1 illustrates the overall pipeline. The details in each stage are as follows.
3.1.1 Concept Categorization
Inspired by recent studies on constructing multicultural evaluation datasets (Liu et al., 2021; Wibowo et al., 2023; Kim et al., 2024), we aim to assess knowledge related to various concepts that Korean natives encounter in daily life. Each concept should not be exclusive to a specific culture but should have some degree of universality. To define the list of such concepts, we refer to semantic concepts provided in the Intercontinental Dictionary Series (IDS) (Key and Comrie, 2015), following Liu et al. (2021). Based on the concepts in the resource and discussions, our dataset is based on top of the following ten concepts: Food, Beverage, Game, Celebrations, Religion, Tool, Clothes, Heritage, Architecture, and Agriculture.
3.1.2 Image Selection
Based on the list of chosen concepts above, Korean native annotators manually collected images from the web. To avoid an excessive number of images containing specific objects, each object is limited to appearing a maximum of two times within any single category. This encourages a broad representation of concepts related to Korean culture. Moreover, following Liu et al. (2021), only images depicting concepts that could physically exist in everyday life were selected. Annotators are encouraged to creatively choose diverse and suitable images by using various web resources. Wikimedia Commons [^1] https://commons.wikimedia.org/wiki are used as the primary sources, and only CC-licensed images are selected.
3.1.3 Question Generation
Question Type
Based on the selected images, we annotate questions in the form of multiple-choice QA. To comprehensively evaluate the understanding of Korean culture, we categorize the questions into two types: visual recognition (Type 1) and reasoning (Type 2). The visual recognition questions ask about basic visual information (e.g., counting, color recognition) or the names of objects appearing in the image. The reasoning questions go beyond simple visual recognition and ask for fine-grained knowledge or reasoning processes related to the given image. For each image, a single Type 1 question and between one and four Type 2 questions were created. This categorization offers two main advantages. First, the Type 1 questions allow for a multifaceted assessment of how well a model understands an image and its embedded concepts. Second, the Type 2 questions ensure that our dataset comprehensively evaluates the understanding of various aspects of Korean culture, rather than merely falling back to simple object recognition.
# of samples 657
- Type 1/ Type 2 237/420
# of unique images 237
# of options 2628
# of unique options 2129
Avg. question length 13.5
- Type 1/ Type 2 10.1/15.5
Avg. option length 1.7
Table 1: Dataset statistics. The length of questions and options denotes the number of words.
Figure 2: Concept distribution of our dataset.
AI-assisted Question Annotation
We create questions and their options (including one correct answer and three distractors) by leveraging a powerful proprietary large VLM (i.e., GPT-4-Turbo). For each concept category, human annotators first write questions and options for at least three images in each category. These manually annotated examples serve as demonstrations for the VLM to generate questions and options along with additional information.
Figure 3: Dataset Examples. We present an image and two questions of different types for each concept category.
In detail, the VLM receives 1) the target image, 2) a set of human-annotated demonstration examples for other images, 3) a detailed guideline for annotation, and 4) an image-wise textual knowledge description. We include a description or general knowledge associated with the given image to enhance the diversity and relevance of the questions. This ensured that the VLM could generate questions grounded in real-world knowledge. One important note is that the guidelines emphasize promoting similarity among all four options for multiple-choice questions, a principle kept in mind while creating the human-annotated examples, inspired by Wang et al. (2023). All textual information is delivered to the VLM through a natural language prompt, and different annotation guidelines are used for generating questions in different types (i.e., visual recognition and reasoning). The text prompts for VLM are presented in Appendix A.
3.1.4 Human Verification
During our preliminary study, we observed that the model often generated inappropriate data samples. For instance, some questions are open-ended or do not induce a single clear answer. In other cases, generated questions are subjective or not sufficiently related to the given image or Korean culture. To ensure the quality of automatically generated datasets, human verification is performed on the generated samples from the VLM. The VLM is guided to generate five question-option sets per image. If none of the five generated samples for a single image meet the criteria, we ask the VLM to regenerate questions and answers. The generated samples deemed appropriate are included in the final dataset with minimal revisions if required. We note that this AI-assisted annotation approach offered several advantages: It significantly accelerated the annotation process and helped minimize unintended human biases. By including human verification on generated questions, we ensure both accuracy and cultural relevance.
3.2 K-Viscuit
Statistics
The detailed statistics of our final benchmark dataset are shown in Table 1. Our dataset comprises a total of 657 examples, with 237 Type 1 questions and 420 Type 2 questions, drawn from 237 unique images from 10 concept categories. The average word count for questions in Type 1 and Type 2 are 10.11 and 15.46, respectively. The average word count for all questions is 13.53 words. Each question is accompanied by four options, resulting in a total of 2,628 options with an average word count of 1.74 words. Fig. 2 illustrates the distribution of selected concepts of samples in our dataset.
Required Knowledge to Solve Questions
To better understand the characteristics of our dataset, we categorize the types of required knowledge to solve questions. We analyzed Type 2 questions using the GPT-4 model following Tong et al. (2024). All Type 2 QA pairs were provided, and the knowledge required to answer each question was summarized. As shown in Fig. 4, diverse cultural elements were considered. Detailed instructions for the categorization can be found in the Appendix B.1.
Figure 4: Required Cultural Knowledge.
Qualitative Examples
Fig. 3 presents sampled images, questions, and options along with their concept category and question type.
4 Experiments
We conduct experiments to evaluate various VLMs on our constructed dataset.
Model All Food Beverage Game Celeb. Religion Tool Clothes Heritage Arch. Agri. InstructBLIP-7B (Dai et al., 2024) 50.84 40.85 42.31 38.46 53.19 40.74 50.67 62.16 51.61 60.55 72.22 instructBLIP-13B (Dai et al., 2024) 55.56 45.77 50.00 46.15 59.57 55.56 54.67 64.86 66.13 60.55 64.81 mPLUG-Owl2-7B (Ye et al., 2023) 48.25 42.25 42.31 30.77 63.83 55.56 48.00 54.05 45.16 49.54 66.67 LLaVA-1.6-7B (Liu et al., 2024) 56.32 43.66 48.08 40.38 57.45 51.85 54.67 67.57 59.68 72.48 72.22 LLaVA-1.6-13B (Liu et al., 2024) 57.08 45.07 53.85 36.54 68.09 40.74 53.33 70.27 66.13 69.72 70.37 InternLM-XC2-7B (Dong et al., 2024) 59.67 50.70 48.08 40.38 65.96 55.56 58.67 64.86 69.35 69.72 75.93 Idefics2-8B (Laurençon et al., 2024) 63.62 51.41 50.00 50.00 74.47 66.67 69.33 75.68 74.19 73.39 62.96 Claude-3-opus (Anthropic, 2024) 70.02 62.68 73.08 59.62 72.34 77.78 74.67 78.38 75.81 67.89 75.93 GPT-4-Turbo (Achiam et al., 2023) 80.82 73.94 80.77 78.85 85.11 85.19 81.33 86.49 85.48 79.82 87.04 Gemini-1.5-Pro (Reid et al., 2024) 81.58 80.28 78.85 71.15 85.11 77.78 82.67 83.78 83.87 84.40 85.19
Table 2: VLMs Evaluation Results. Celeb., Arch., and Agri. denote Celebration, Architecture, and Agriculture. The highest and the second highest scores in each column are highlighted in bold and underlined.
4.1 Models
The following open-source VLMs are used for experiments: InstructBLIP-7B/13B (Dai et al., 2024), LLaVA-v1.6-7B/13B (Liu et al., 2024), mPLUG-Owl2-7B (Ye et al., 2023), InternLM-XComposer2-VL-7B (Dong et al., 2024), and Idefics2-8B Laurençon et al. (2024). We also use the following proprietary models: Claude-3-opus (Anthropic, 2024), GPT-4-Turbo (Achiam et al., 2023), and Gemini-1.5-Pro (Reid et al., 2024). All models are evaluated in the conventional multiple-choice setup, where a model is prompted to choose its answer from one of four options. The input text is constructed by concatenating (1) a question, (2) each option with option letters in alphabetical order, and (3) the instruction about output format (i.e., "Answer with the option’s letter from the given choices directly."). We use accuracy as an evaluation metric.
Model Type 1 Type 2 Total
InstructBLIP-7B 45.57 53.81 50.84
InstructBLIP-13B 51.48 57.86 55.56
mPLUG-Owl2-7B 43.04 51.19 48.25
LLaVA-1.6-7B 50.21 59.76 56.32
LLaVA-1.6-13B 54.01 58.81 57.08
InternLM-XC2-7B 56.12 61.67 59.67
Idefics2-8B 63.71 63.57 63.62
Claude-3-opus 69.62 80.24 70.02
GPT-4-Turbo 78.90 81.90 80.82
Gemini-1.5-Pro 83.97 70.24 81.58
Table 3: Results on Different Question Types. Type 1 and Type 2 denote question types in visual recognition and visual reasoning, respectively.
4.2 Results
Table 2 demonstrates the evaluation results of different VLMs on our dataset. We first observe that proprietary models usually show higher accuracies. The Gemini-1.5-Pro and GPT-4-Turbo achieve the highest and second-highest scores, respectively. Among the open-sourced models, Idefics2 and InternLM-XComposer2-VL usually perform better than other models. The accuracy of models in different question types is shown in Table 3. We observe that most models show higher accuracy in Type 2 questions compared to Type 1 questions, except for Gemini-1.5-Pro and Idefics2-8B. Regarding these trends, we suspect visual recognition with diverse cultural contexts poses inherent challenges for VLMs. Our dataset enables the evaluation and identification of such aspects where models can be improved.
Figure 5: Qualitative examples of selected VLMs (i.e., LLaVA-1.6-13B, InternLM-XC2-7B, GPT-4-Turbo, and Gemini-1.5-Pro) on sampled questions. The answer option is highlighted in the underline. The correct and incorrect choice of models are highlighted in green and red.
4.3 Analyses
Asking the VLM in Korean
In this work, we primarily focused on evaluating the VLM’s understanding of Korean culture and intentionally excluded their understanding of the Korean language. Thus, we designed the dataset to focus on the VLM’s multiculturality, independent of its multilingualism. However, cultural information about certain groups often exists in the language that persons in the group frequently use. In other words, VLMs trained in multilingual corpora might have learned about Korean culture through texts written in Korean. Therefore, we probe whether asking questions in Korean could improve the performance of VLMs on our dataset. To translate our dataset into Korean, we use proprietary unimodal LM (gpt-3.5-turbo) as a machine translation system. For each sample in our dataset, the LM translates questions and four options into Korean. Three proprietary VLMs that can receive Korean text (i.e., GPT-4-Turbo, Claude-3-opus, and Gemini-1.5-Pro) are used for experiments.
Evaluation results with different input languages are presented in Table 4. We observe that solely providing translated texts in Korean to VLMs does not contribute to model performance. When English texts are also given to models, Gemini-1.5-Pro shows increased performance (i.e., 81.58 to 83.41). GPT-4-Turbo and Claude-3-opus usually do not take the benefits of using Korean texts.
Model en ko en + ko
Claude-3-Opus 70.02 65.30 70.32
GPT-4-Turbo 80.82 75.34 80.37
Gemini-1.5-Pro 81.58 80.82 83.41
Table 4: Results with Different Input Languages. ko is machine-translated texts. For en+ko, we provide questions and options in both languages to models.
Human Evaluation on the Benchmark
To evaluate how well people from different backgrounds perform on K-Viscuit, we conducted human evaluation. We selected a subset of our dataset by randomly sampling 25 images, each with one type 1 and one type 2 question, totaling 50 questions. This test was administered to 20 Koreans and 14 Non-Koreans. As depicted in Fig. 6, the results showed that participants with better knowledge of Korean culture achieved higher accuracy. However, even among Koreans, knowledge gaps exist both between individuals and within a single person’s knowledge across different domains, such as history. Proprietary VLMs can match or exceed human performance, indicating that utilizing VLMs for generating diverse, culturally nuanced questions is more effective than relying solely on individual human annotators. Details of the user study are provided in the Appendix B.2.
Figure 6: Human Evaluation. Accuracy comparison between Koreans and non-Koreans on a K-Viscuit subset, measured across different proficiency levels.
Qualitative Results
Fig. 5 presents the prediction of selected models on sampled examples. In the first example, all models fail to correctly answer the question about asking detailed game rules. In the second case, two proprietary models are wrong while open-source models make correct answers. The third problem requires the model to recognize the structure in the image as Cheomseongdae and infer that it serves a similar function to the Griffith Observatory in the United States. Both proprietary models make correct answers to this example. In the final example, all models successfully identify the correct answer about agriculture.
Figure 7: Qualitative results of a VLM with multi-choice and generative VQA setups. LLaVA-1.6-13B is used for the analysis. The correct and incorrect model generations are manually highlighted by the authors.
5 Discussion
We discuss several components and future directions to build VLMs grounded on diverse cultures.
5.1 Evaluation beyond Multiple-choice VQA
Our dataset is designed as a multiple-choice VQA task, where VLMs are required to choose one of the given options. While this classification setup is advantageous for measuring model performance through accuracy, it may not fully capture the models’ understanding of Korean culture. For instance, VLMs may correctly select an answer from the given options but fail in a generative VQA setup where they must generate a free-form answer. An example of this is shown in Fig. 7 (left), where the VLM correctly identifies an object from the choices but hallucinates it as “stones or bricks” when generating an answer without options. In contrast, the right example in Fig. 7 shows the VLM accurately identifying the object as “bibimbap” in both multiple-choice and generative setups. Our future research aims to propose various evaluation methods to assess whether VLMs genuinely possess the knowledge to answer questions accurately.
5.2 Augmenting VLMs with External Knowledge
In previous experiments, we observed significant performance gaps in models for questions that require cultural knowledge. To address these gaps, two methodologies can be considered: fine-tuning open-source models and leveraging external knowledge. While fine-tuning is limited to open-source models, leveraging external knowledge can be applied to both open-source and proprietary models. Therefore, we focused on the latter approach.
We augmented the models with relevant documents for each test image from the Food concept. Our retrieval method involved generating captions using GPT-4-Turbo, embedding these captions via text-embedding-3-large to build queries, and retrieving Top-1 document from 152 Wikipedia pages related to the objects mentioned in the Type 1 options. The distractors in K-Viscuit were designed to be highly similar to the correct answers. This creates a challenging retrieval setting with numerous hard negatives, aiming to mimic real-world conditions. We aim to assess whether the retrieval method remains effective even in such a setting by observing the performance changes when cultural knowledge is added. The results show that providing retrieved documents can improve model performance, as indicated by higher scores in the Retrieved setting compared to None. The Oracle setting, which uses a curated relevant document, serves as an upper bound. Overall, augmenting VLMs with external knowledge via document retrieval demonstrates the potential to bridge the performance gap in VQA tasks requiring cultural context. We provide the details in the Appendix D.
Model None Retrieved Oracle
LLaVA-1.6-7B 43.66 68.31 78.87
LLaVA-1.6-13B 45.77 64.08 80.28
InternLM-XC2-7B 50.70 68.31 82.39
Idefics2-8B 51.41 67.61 82.39
Claude-3-opus 62.68 70.42 87.32
GPT-4-Turbo 73.94 78.17 88.73
Gemini-1.5-Pro 80.28 78.17 90.85
Table 5: Retrieval-augmented Generation Results.
6 Conclusion
In this work, we proposed a semi-automated pipeline to construct culturally aware benchmarks for vision-language models, featuring multi-choice QA with two types of questions that require visual recognition or reasoning. This pipeline combines human-VLM collaboration, enhancing efficiency and diversity in generating culturally relevant QAs. Our approach, first used to create the K-Viscuit dataset, can be adapted to other cultures.
Our evaluation with K-Viscuit identified a notable performance gap between open-source and proprietary VLMs in understanding Korean culture, highlighting the need for improvement in open-source models. Further analyses demonstrated the impact of in-depth cultural understanding on VQA results. Finally, we enhanced VLM performance through knowledge augmentation and explored open-ended answer generation beyond the multiple-choice setting. Our work stresses the importance of cultural diversity in evaluation, aiming for more inclusive and accurate VLMs in the future.
Limitations
Our current framework requires a manual selection of images that match specified concepts, preventing fully automated dataset generation. These human efforts can be alleviated by multimodal retrieval modules to some extent. To this end, it should be predetermined whether current multimodal encoders can sufficiently understand culturally nuanced images and texts. Enhancing retrieval models to better understand and match cultural contexts remains our exciting future work. The manual verification of automatically generated questions also can be a considerable burden. Developing a quality estimation module for generated questions could assist in this process by reducing the workload on human annotators.
Appendix A Dataset Construction Details
Guidelines to GPT-4-Turbo for Dataset Annotation
The detailed prompts for the annotation with GPT-4-Turbo are presented in Table 6 and Table 7.
Prompt for Type 1 annotations:
[System Prompt]
You are a helpful Korean annotator to make visual question answering datasets.
[User Prompt]
Given an image, generate a question asking for the name of the main object or the main activity that people are engaged in, and generate one correct option (answer) and three wrong options (distractors).
Detailed guidelines are as follows: 1. The objects shown in the image is called ‘‘{object_name}’’ in Korean. You can include this word into your correct options after translation into English. 2. All options should be written in up to 5 words. 3. Please struggle to make creative or challenging distractors so that they are not easily distinguished from the answer. 4. Distractors should seem similar to the correct answer and related to the category of the main object in image (e.g., Hanok - Agungi, Sarangchae, Anchae, Daecheongmaru). Distractors are better when they have similar color, shape, or texture with the answer. 5. Separate each distractor with “;” symbol. 6. Don’t make any explanation. 7. Distractors should be culturally related to the image. 8. All the options should be either transliterated or translated. Never mix transliteration and translation. 9. Don’t be too specific (Avoid using a proper name: instead use [University building] instead of [Ewha Campus Complex]) You can refer to below examples that are annotated for other images.
Question: What is the name of this place shown in the image?
Answer: Sarangchae Distractors: Anchae ; Sadang ; Daecheongmaru Question: What is the name of the structure seen in the image? Answer: Ondol Distractors: Agungi ; Jangdokdae ; Buttumak Question: What is the name of this building shown in the image? Answer: Gosiwon Distractors: Officetel ; Apartment ; Share house Please make four options (single answer and three distractors).
Table 6: A prompt of GPT-4-Turbo used in the annotation of Type 1 questions (i.e., visual recognition) in the Architecture category.
Prompt for Type 2 annotations:
[System Prompt]
You are a helpful Korean annotator to make visual question answering datasets.
[User Prompt]
Please ask 5 questions and their options about the image. Here are the guidelines to follow for writing.
Detailed guidelines are as follows:
1. The objects shown in the image is ‘‘{object_name}’’. Don’t include this word in your questions.
[…]
You can refer to below examples that are annotated for other images. Question: Seen in the image, what traditional Korean heated floor system is associated with the heat source from this feature? Answer: Ondol Distractors: Daecheongmaru ; Anchae ; Buttumak Question: In the image, what kind of Korean roof finishing is visible, known for its multicolored patterns? Answer: Dancheong Distractors: Seoggarae ; Cheoma ; Maru Question: Which mode of transportation is commonly used by tourists to ascend the mountain where the tower is located? Answer: Cable Car Distractors: Bus ; Bicycle ; Funicular Railway Please make four options (single answer and three distractors).
Table 7: A prompt of GPT-4-Turbo used in the annotation of Type 2 questions (i.e., visual reasoning) in the Architecture category.
Appendix B Dataset Analyses Details
B.1 Required Knowledge Analysis
We analyzed how diverse the cultural knowledge required by the questions in our K-Viscuit dataset is. To this end, we used the following prompt to obtain responses from the GPT-4 model. We delivered all Type 2 samples (including both the questions and the options) to the model by concatenating them into a single string.
Prompt for required knowledge analysis:
I created a multiple-choice quiz with four options per question, based on images related to Korean culture. Each question is designed to assess the understanding of one or more cultural elements. Please analyze which cultural element each question aims to measure and provide an overall summary. ‘‘{TYPE 2 samples}’’
Table 8: Prompt for required knowledge analysis.
B.2 Human Evaluation Details
We randomly selected images according to the proportion of each category to create the questionnaire for the human evaluation. If there were multiple Type 2 questions for a single image, we sampled them randomly. The number of selected images per category is as follows: Food (4), Beverage (2), Game (2), Celebrations (2), Religion (2), Tool (3), Clothes (2), Heritage (2), Architecture (4), and Agriculture (2).
We released the survey on the Amazon MTurk platform, where non-Koreans with a relatively limited understanding of Korean culture were asked to complete the K-Viscuit questions within 20 minutes for a compensation of \$5. The survey on the MTurk platform resulted in a demographic composed entirely of Americans. Their self-assessed proficiency levels were: Very familiar (35.7%), Somewhat familiar (50%), Slightly familiar (14.3%), and Not familiar at all (0%). For Koreans, we administered the survey to 20 graduate students in their mid-to-late twenties. We received feedback that Koreans had the most difficulty with questions related to history.
Appendix C VLM Evaluation Details
Model Implementation Details
We present further implementation details in VLMs used in our experiments. All open-source VLMs are implemented with the Transformers framework (Wolf et al., 2020), and the checkpoints are downloaded from Huggingface Hub [^2] https://huggingface.co/models. For proprietary models, gpt-4-turbo-2024-04-09, gemini-1.5-pro, and claude-3-opus -20240229 are used. The text prompt used for proprietary models is presented in Table 9.
Inference prompt for proprietary VLMs:
[System Prompt]
You will be given an image taken in Korea and a 4-way multiple-choice question. Answer the question based on the given image and your knowledge about Korean culture.
[User Prompt]
Question: ‘‘{question}’’
Options:
a. ‘‘{option_a}’’
b. ‘‘{option_b}’’
c. ‘‘{option_c}’’
d. ‘‘{option_d}’’
Make sure to respond with the option’s letter: ‘a.’, ‘b.’, ‘c.’, or ‘d.’. Do not make any additional explanation.
Table 9: Inference prompt for proprietary VLMs.
Appendix D Retrieval Methodology Details
For external knowledge retrieval, we generated image captions using GPT-4 with the prompt shown in Table 10.
Prompt for generating image captions:
You are an AI language model specializing in identifying and describing Korean food from photographs. When given a photograph of Korean food, your task is to accurately describe the food based on its visual characteristics and visible ingredients. Your description should include the name of the dish, main ingredients, common accompaniments, and notable features that help identify the food. Be detailed yet concise, providing clear and helpful information to those trying to understand Korean cuisine. Ensure that your description is within 150 words.
Table 10: Prompt for generating image captions.

### v2

@@
 Abstract
-To create culturally inclusive vision-language models (VLMs), the foremost requirement is developing a test benchmark that can diagnose the models’ ability to respond to questions reflecting cultural elements.
-This paper addresses the necessity for such benchmarks, noting that existing research has relied on human annotators’ manual efforts, which impedes diversity and efficiency.
-We propose a semi-automated pipeline for constructing cultural VLM benchmarks to enhance diversity and efficiency.
-This pipeline leverages human-VLM collaboration, where VLMs generate questions based on guidelines, human-annotated examples, and image-wise relevant knowledge, which are then reviewed by native speakers for quality and cultural relevance.
-The effectiveness of our adaptable pipeline is demonstrated through a specific application: creating a dataset tailored to Korean culture, dubbed K-Viscuit.
-The resulting benchmark features two types of questions: Type 1 questions measure visual recognition abilities, while Type 2 assess fine-grained visual reasoning skills.
-This ensures a thorough diagnosis of VLM models across various aspects.
-Our evaluation using K-Viscuit revealed that open-source models notably lag behind proprietary models in understanding Korean culture, highlighting areas for improvement.
-We provided diverse analyses of VLM performance across different cultural aspects.
-Besides, we explored the potential of incorporating external knowledge retrieval to enhance the generation process, suggesting future directions for improving cultural interpretation ability of VLMs.
-Our dataset and code will be made publicly available.
+To create culturally inclusive vision-language models (VLMs), developing a benchmark that tests their ability to address culturally relevant questions is essential.
+Existing approaches typically rely on human annotators, making the process labor-intensive and creating a cognitive burden in generating diverse questions.
+To address this, we propose a semi-automated framework for constructing cultural VLM benchmarks, specifically targeting multiple-choice QA.
+This framework combines human-VLM collaboration, where VLMs generate questions based on guidelines, a small set of annotated examples, and relevant knowledge, followed by a verification process by native speakers.
+We demonstrate the effectiveness of this framework through the creation of K-Viscuit, a dataset focused on Korean culture.
+Our experiments on this dataset reveal that open-source models lag behind proprietary ones in understanding Korean culture, highlighting key areas for improvement.
+We also present a series of further analyses, including human evaluation, augmenting VLMs with external knowledge, and the evaluation beyond multiple-choice QA.
+[^1] Our dataset is available at https://huggingface.co/datasets/ddehun/k-viscuit.
 Evaluating Visual and Cultural Interpretation:
@@
 1 Introduction
-The utility of vision-language models (VLMs) in bridging visual and textual data is well-recognized, offering advancements in applications such as image captioning and optical character recognition.
-However, most VLMs are predominantly trained on data from Western languages and cultures (Lin et al., 2014; Young et al., 2014; Antol et al., 2015), limiting their effectiveness in non-Western contexts (Liu et al., 2021; Yin et al., 2023, 2021; Romero et al., 2024).
-Cultural nuances significantly influence the interpretation of visual and textual information, impacting the accuracy and relevance of VLM outputs.
-Consequently, building VLMs that understand diverse cultural backgrounds is essential for enhancing inclusivity and ensuring fair performance across different demographics.
-To develop VLMs that effectively understand the cultural contexts of specific countries, establishing a wide-ranging test benchmark is a critical first step.
-Such a benchmark is essential for diagnosing the model’s ability to accurately interpret and respond to culturally specific inputs.
-Notably, previous research has endeavored to create VLM test benchmarks for particular countries (Liu et al., 2021; Yin et al., 2021; Romero et al., 2024).
-Figure 1: Pipeline Overview.
-However, these approaches have primarily relied on manually crafted questions, which present two major limitations.
-First, their process is time-consuming and labor-intensive, increasing the cost of the research and making it difficult to extend and apply to the cultural contexts of other countries.
-Additionally, the diversity of questions can be limited due to cognitive fixation, where humans find it challenging to generate a wide range of varied questions after settling on an initial idea (Ramos, 2020).
-This lack of diversity can hinder the accurate diagnosis of the model’s ability to generalize across different types of culturally nuanced inputs.
-To address these limitations, we propose a semi-automated pipeline for constructing cultural VLM benchmarks that enhance the efficiency and diversity of culture-relevant visual question and answer generation as illustrated in Fig.
+Recent advances in vision-language models (VLMs) have demonstrated remarkable capabilities in tasks ranging from image captioning to visual question answering.
+However, these models are predominantly trained on Western-centric datasets (Lin et al., 2014; Young et al., 2014; Antol et al., 2015), leading to significant performance disparities when applied to non-Western contexts (Liu et al., 2021; Yin et al., 2021, 2023; Romero et al., 2024).
+This cultural bias is particularly problematic as visual interpretation often depends heavily on cultural context, necessitating the development of more culturally aware VLMs.
+Several benchmarks have been proposed to evaluate cultural understanding in VLMs (Liu et al., 2021; Yin et al., 2021; Romero et al., 2024; Nayak et al., 2024).
+These approaches primarily rely on manual question generation, which, while valuable, faces certain practical challenges.
+The manual process can be time-consuming and resource-intensive when scaling to new cultural contexts.
+Additionally, as noted in cognitive science research (Ramos, 2020), human annotators may experience cognitive fixation, potentially limiting the diversity of generated questions.
+These practical considerations motivate the need for more efficient benchmark construction approaches.
+Figure 1: Framework Overview.
+Inspired by recent successes in human-LLM collaborative data generation (Liu et al., 2022; Bartolo et al., 2022; Kamalloo et al., 2023), we propose a semi-automated framework for constructing cultural VLM benchmarks that enhances both the efficiency and diversity of culture-relevant visual question and answer generation, as shown in Fig.
 1.
-This pipeline incorporates human-VLM collaboration, where the VLM generates and recommends questions and answers leveraging carefully crafted guidelines, a small set of human-annotated examples, and image-wise relevant knowledge.
-These recommendations are then assessed and approved by native speakers, ensuring high quality and cultural relevance.
-In this work, we applied the proposed pipeline to create a dataset for Korean culture, named K-Viscuit (Korean Visual and Cultural Interpretation Test).
-Nevertheless, our pipeline can be adapted and applied to other cultures analogously.
-The resulting benchmark dataset created using our proposed pipeline has two notable features.
-Firstly, it has the ability to measure nuanced cultural understanding through two types of assessments: Type 1, which measures visual recognition capability, and Type 2, which measures visual reasoning capability.
-Secondly, the multiple-choice questions in the dataset include four options, with distractors generated following carefully designed instructions to ensure high similarity among the options.
-This high similarity between options prevents models from using shortcuts to determine the correct answer, ensuring a more robust and comprehensive evaluation of cultural understanding in VLMs.
-These features underscore the effectiveness and thoroughness of the resulting benchmark dataset in addressing cultural nuances.
-Using our benchmark data, we evaluated the performance of various open-source and proprietary VLMs.
-Our experimental results indicate that open-source models significantly lag behind proprietary models in the context of Korean culture.
-This assessment highlights key improvement areas and sets development goals for open-source models, with proprietary models as the reference.
+Our framework incorporates human-VLM collaboration, where the VLM generates and recommends questions and answers based on carefully crafted guidelines, a small set of human-annotated examples, and image-specific knowledge.
+Native speakers then verify these recommended questions to ensure quality and cultural relevance.
+Using this framework, we develop K-Viscuit (Korean Visual and Cultural Interpretation Test), a benchmark dataset for Korean culture that can be adapted for other cultural contexts.
+K-Viscuit features two distinct evaluation types: visual recognition and visual reasoning.
+Also, the benchmark employs carefully designed multiple-choice questions with highly similar distractors to prevent models from exploiting superficial patterns.
+Our evaluation with K-Viscuit reveals a significant performance gap between open-source and proprietary VLMs in understanding Korean culture.
+We provide insights into the current limitations and potential improvements in VLMs’ cultural understanding capabilities through detailed analyses, which include human evaluation, external knowledge integration, and extended evaluation beyond a multi-choice question-answering setup.
 Our contributions are summarized as follows:
-  • We propose a pipeline for building benchmarks to diagnose VLMs’ understanding of various cultures. It enables efficient creation of culturally diverse datasets through human-VLM cooperation.
-  • Using our pipeline, we constructed the K-Viscuit Korean Culture VQA benchmark.
-  • We utilized K-Viscuit to diagnose and analyze the performance of existing open-source and proprietary VLMs.
+  • We propose a semi-automated framework for constructing benchmarks to evaluate the cultural understanding capabilities of VLMs.
+  • We develop K-Viscuit, a Korean culture-focused VQA benchmark using our proposed framework.
+  • We present comprehensive experimental results and analyses of both open-source and proprietary VLMs evaluated on K-Viscuit.
 2 Related Work
-Recent research has made significant strides in developing benchmarks to assess the ability of AI models, particularly VLMs and Large Language Models (LLMs), to understand and respond appropriately to various cultural contexts.
-These efforts are crucial given that many existing models are trained predominantly on Western-centric datasets (Young et al., 2014; Lin et al., 2014; Antol et al., 2015), which constrains their performance in non-Western settings.
-On the side of LLMs, Kim et al.
-(2024) introduced a benchmark dataset, CLIcK, specifically for assessing cultural and linguistic intelligence in Korean language models.
-This dataset includes QA pairs designed to test the cultural knowledge of Korean language models.
-Wibowo et al.
-(2023) created COPAL-ID, a dataset that captures Indonesian cultural nuances for text-based commonsense reasoning, highlighting the need for culturally aware benchmarks.
-VLMs require consideration of both image and text inputs to effectively reflect cultural contexts, as cultural nuances can arise from either source.
+Recent research has made significant strides in developing benchmarks to assess AI models’ cultural understanding capabilities.
+These efforts are particularly important as many existing models, including VLMs and Large Language Models (LLMs), are trained predominantly on Western-centric datasets (Young et al., 2014; Lin et al., 2014; Antol et al., 2015), limiting their effectiveness in non-Western contexts.
+For LLMs, several notable cultural benchmarks have emerged: Kim et al.
+(2024) introduced CLIcK, a benchmark for evaluating Korean language models’ cultural knowledge through carefully designed QA pairs, while Wibowo et al.
+(2023) developed COPAL-ID to capture Indonesian cultural nuances in text-based commonsense reasoning.
+In the multimodal domain, VLMs require consideration of both visual and textual inputs to effectively reflect cultural contexts.
 Liu et al.
-(2021) proposed the MaRVL dataset, a multilingual visually grounded reasoning dataset that includes five different languages and cultures to diagnose inherent language and cultural biases in pre-trained models.
-Their work highlights the need for models to seamlessly interpret and integrate visual and textual information across diverse cultures, paving the way for more inclusive and globally aware AI systems.
-Yin et al.
-(2021) built the GD-VCR dataset, which tests visual commonsense reasoning influenced by geographical and cultural contexts.
-Romero et al.
-(2024) presented the CVQA, a multilingual visual question-answering (VQA) benchmark that incorporates various languages and cultures, providing a comprehensive tool for evaluating VLMs in diverse cultural contexts.
-Previous efforts to create cultural benchmarks for VLMs, however, have heavily relied on human annotators, which can limit both the diversity and efficiency of dataset creation (Liu et al., 2021; Yin et al., 2021; Ramaswamy et al., 2024; Romero et al., 2024).
-The efficiency and diversity of dataset creation can be significantly enhanced by leveraging AI models, particularly when combined with human collaboration.
-This approach not only improves the speed and scope of dataset generation but also maintains high-quality standards.
+(2021) addressed this challenge with MaRVL, a multilingual visually-grounded reasoning dataset spanning five languages and cultures, demonstrating the importance of cross-cultural visual-linguistic understanding.
+Building on this foundation, Yin et al.
+(2021) introduced GD-VCR to evaluate geographical and cultural aspects of visual commonsense reasoning, while Romero et al.
+(2024) developed CVQA, a comprehensive multilingual VQA benchmark for assessing VLMs across diverse cultural contexts.
+While these cultural benchmarks have provided valuable insights, their reliance on manual annotation can constrain both the diversity and efficiency of dataset creation (Liu et al., 2021; Yin et al., 2021; Ramaswamy et al., 2024; Romero et al., 2024).
+Recent work has demonstrated the potential of AI-assisted dataset generation when combined with human expertise.
 Liu et al.
-(2022) demonstrated this in their study on building a natural language inference dataset, where they combined the generative capabilities of language models with human review.
-Analogous to how LLMs are heavily utilized in the creation of datasets for language-only tasks (Taori et al., 2023; Kim et al., 2023; Dubois et al., 2024; Kim et al., 2024), leveraging the inherent knowledge of AI models in building multimodal cultural benchmarks can yield diverse and culturally nuanced examples.
-In this work, by harnessing AI models, we have significantly enhanced the diversity and richness of the dataset, ensuring it better reflects the wide array of cultural contexts we aim to represent.
-3 Data Construction Pipeline
-We present our human-AI collaborated dataset construction pipeline to diagnose the understanding of VLMs in a target cultural field.
-In this work, we focus on Korean culture as a target cultural field.
-First, we provide an overview (section 3.1) and details of the pipeline.
-Then, we present the analysis of the resulting dataset, dubbed K-Viscuit (Korean Visual and Cultural Interpretation Test) (section 3.2).
-3.1 Pipeline Overview
-The pipeline is designed to comprise a multiple-choice visual question answering (VQA) task, where each evaluation sample consists of an image, a question, and four options with one correct choice.
-Native Korean speakers are involved in the dataset construction, and a powerful proprietary VLM is employed to mitigate unintended human biases, such as cognitive fixation (Ramos, 2020), and encourage an effective annotation process.
-The generated samples are designed to cover various aspects of a target culture derived from daily life and necessitate multimodal reasoning to interpret both the image and the question accurately.
-All text in the dataset is written in English to isolate the evaluation of multicultural comprehension from multilingual aspects.
+(2022) successfully applied this approach to natural language inference tasks, and similar strategies have been widely adopted in creating language-only datasets (Taori et al., 2023; Kim et al., 2023; Dubois et al., 2024; Kim et al., 2024).
+Our work extends this paradigm to multimodal cultural benchmarks, leveraging AI models to enhance dataset diversity while maintaining high quality through human verification.
+Figure 2: Dataset Examples. We present an image and two questions of different types for each concept category.
+3 Data Construction Framework
+We present our human-AI collaborative framework for constructing datasets to evaluate VLMs’ understanding of specific cultural domains.
+In this work, we focus on Korean culture as our target domain.
+First, we provide an overview (section 3.1) and implementation details.
+Then, we present the analysis of our resulting dataset, K-Viscuit (Korean Visual and Cultural Interpretation Test) (section 3.2).
+3.1 Framework Overview
+Our framework is designed to create a multiple-choice visual question answering (VQA) task, where each evaluation sample consists of an image, a question, and four options with one correct answer.
+Native Korean speakers participate in the dataset construction process, while a powerful proprietary VLM is employed to mitigate unintended human biases, such as cognitive fixation (Ramos, 2020), and streamline the annotation process.
+The generated samples cover various aspects of the target culture derived from daily life and require multimodal reasoning to interpret both visual and textual information accurately.
 The dataset construction consists of four stages: 1) concept selection, 2) image selection, 3) question and options annotation, and 4) human verification.
 Fig.
-1 illustrates the overall pipeline.
-The details in each stage are as follows.
+1 illustrates the overall framework.
 3.1.1 Concept Categorization
-Inspired by recent studies on constructing multicultural evaluation datasets (Liu et al., 2021; Wibowo et al., 2023; Kim et al., 2024), we aim to assess knowledge related to various concepts that Korean natives encounter in daily life.
-Each concept should not be exclusive to a specific culture but should have some degree of universality.
-To define the list of such concepts, we refer to semantic concepts provided in the Intercontinental Dictionary Series (IDS) (Key and Comrie, 2015), following Liu et al.
-(2021).
-Based on the concepts in the resource and discussions, our dataset is based on top of the following ten concepts: Food, Beverage, Game, Celebrations, Religion, Tool, Clothes, Heritage, Architecture, and Agriculture.
+Inspired by recent studies on multicultural evaluation datasets (Liu et al., 2021; Wibowo et al., 2023; Kim et al., 2024), we aim to assess knowledge of various concepts encountered in daily life by Korean natives.
+While each concept should have some degree of universality, its manifestation often varies across cultures.
+Following Liu et al.
+(2021), we reference semantic concepts from the Intercontinental Dictionary Series (IDS) (Key and Comrie, 2015) to define our concept list.
+Our dataset encompasses ten core concepts: Food, Beverage, Game, Celebrations, Religion, Tool, Clothes, Heritage, Architecture, and Agriculture.
 3.1.2 Image Selection
-Based on the list of chosen concepts above, Korean native annotators manually collected images from the web.
-To avoid an excessive number of images containing specific objects, each object is limited to appearing a maximum of two times within any single category.
-This encourages a broad representation of concepts related to Korean culture.
-Moreover, following Liu et al.
-(2021), only images depicting concepts that could physically exist in everyday life were selected.
-Annotators are encouraged to creatively choose diverse and suitable images by using various web resources.
-Wikimedia Commons [^1] https://commons.wikimedia.org/wiki are used as the primary sources, and only CC-licensed images are selected.
+Korean native annotators collected web images corresponding to the selected concepts.
+To ensure diverse representation, we limited each specific object to appearing no more than twice within any single category.
+Following Liu et al.
+(2021), we selected only images depicting concepts that could physically exist in everyday life.
+Annotators were encouraged to source diverse and suitable images from various web resources.
+Wikimedia Commons [^2] https://commons.wikimedia.org/wiki served as the primary source, and only CC-licensed images were selected.
 3.1.3 Question Generation
 Question Type
-Based on the selected images, we annotate questions in the form of multiple-choice QA.
-To comprehensively evaluate the understanding of Korean culture, we categorize the questions into two types: visual recognition (Type 1) and reasoning (Type 2).
-The visual recognition questions ask about basic visual information (e.g., counting, color recognition) or the names of objects appearing in the image.
-The reasoning questions go beyond simple visual recognition and ask for fine-grained knowledge or reasoning processes related to the given image.
-For each image, a single Type 1 question and between one and four Type 2 questions were created.
-This categorization offers two main advantages.
-First, the Type 1 questions allow for a multifaceted assessment of how well a model understands an image and its embedded concepts.
-Second, the Type 2 questions ensure that our dataset comprehensively evaluates the understanding of various aspects of Korean culture, rather than merely falling back to simple object recognition.
+Based on the selected images, we annotate questions in multiple-choice QA format.
+To comprehensively evaluate understanding of Korean culture, we categorize questions into two types: visual recognition (Type 1) and reasoning (Type 2).
+Visual recognition questions assess basic visual information such as object identification, while reasoning questions require fine-grained cultural knowledge or deeper reasoning processes related to the image.
+For each image, we created one Type 1 question and between one to four Type 2 questions.
+This categorization offers two key advantages: First, Type 1 questions enable assessment of a model’s basic visual understanding of culturally embedded concepts.
+Second, Type 2 questions comprehensively evaluate cultural understanding beyond simple object recognition.
 # of samples 657
@@
 Table 1: Dataset statistics. The length of questions and options denotes the number of words.
+AI-assisted Question Annotation
+We create questions and their options (one correct answer and three distractors) by leveraging a powerful proprietary VLM (GPT-4-Turbo).
+For each concept category, human annotators first create exemplar questions and options for at least three images.
+These manually annotated examples serve as demonstrations for the VLM to generate additional questions and options.
+Specifically, the VLM receives: 1) the target image, 2) human-annotated demonstration examples, 3) detailed annotation guidelines, and 4) image-specific knowledge descriptions.
+We include relevant contextual knowledge for each image to enhance question diversity and relevance, ensuring VLM-generated questions are grounded in real-world understanding.
+Notably, following Wang et al.
+(2023), our guidelines emphasize maintaining high similarity among all four multiple-choice options, a principle also reflected in human-annotated examples.
+All information is provided to the VLM through natural language prompts, with distinct annotation guidelines for visual recognition and reasoning questions.
+The detailed prompts are presented in Appendix A.
+It should be noted that all text in the dataset is written in English to isolate the evaluation of multicultural comprehension from multilingual aspects.
+However, as culture and language are not entirely orthogonal (SUSAN, 1996; Kramsch, 2014), completely separating them can be challenging.
+For instance, we observed that certain Korean terms lack exact English equivalents.
+In such cases, we romanize these Korean terms following standard transliteration rules.
 Figure 3: Concept distribution of our dataset.
-AI-assisted Question Annotation
-We create questions and their options (including one correct answer and three distractors) by leveraging a powerful proprietary large VLM (i.e., GPT-4-Turbo).
-For each concept category, human annotators first write questions and options for at least three images in each category.
-These manually annotated examples serve as demonstrations for the VLM to generate questions and options along with additional information.
-Figure 3: Dataset Examples. We present an image and two questions of different types for each concept category.
-In detail, the VLM receives 1) the target image, 2) a set of human-annotated demonstration examples for other images, 3) a detailed guideline for annotation, and 4) an image-wise textual knowledge description.
-We include a description or general knowledge associated with the given image to enhance the diversity and relevance of the questions.
-This ensured that the VLM could generate questions grounded in real-world knowledge.
-One important note is that the guidelines emphasize promoting similarity among all four options for multiple-choice questions, a principle kept in mind while creating the human-annotated examples, inspired by Wang et al.
-(2023).
-All textual information is delivered to the VLM through a natural language prompt, and different annotation guidelines are used for generating questions in different types (i.e., visual recognition and reasoning).
-The text prompts for VLM are presented in Appendix A.
-3.1.4 Human Verification
-During our preliminary study, we observed that the model often generated inappropriate data samples.
-For instance, some questions are open-ended or do not induce a single clear answer.
-In other cases, generated questions are subjective or not sufficiently related to the given image or Korean culture.
-To ensure the quality of automatically generated datasets, human verification is performed on the generated samples from the VLM.
-The VLM is guided to generate five question-option sets per image.
-If none of the five generated samples for a single image meet the criteria, we ask the VLM to regenerate questions and answers.
-The generated samples deemed appropriate are included in the final dataset with minimal revisions if required.
-We note that this AI-assisted annotation approach offered several advantages: It significantly accelerated the annotation process and helped minimize unintended human biases.
-By including human verification on generated questions, we ensure both accuracy and cultural relevance.
-3.2 K-Viscuit
-Statistics
-The detailed statistics of our final benchmark dataset are shown in Table 1.
-Our dataset comprises a total of 657 examples, with 237 Type 1 questions and 420 Type 2 questions, drawn from 237 unique images from 10 concept categories.
-The average word count for questions in Type 1 and Type 2 are 10.11 and 15.46, respectively.
-The average word count for all questions is 13.53 words.
-Each question is accompanied by four options, resulting in a total of 2,628 options with an average word count of 1.74 words.
-Fig.
-2 illustrates the distribution of selected concepts of samples in our dataset.
-Required Knowledge to Solve Questions
-To better understand the characteristics of our dataset, we categorize the types of required knowledge to solve questions.
-We analyzed Type 2 questions using the GPT-4 model following Tong et al.
-(2024).
-All Type 2 QA pairs were provided, and the knowledge required to answer each question was summarized.
-As shown in Fig.
-4, diverse cultural elements were considered.
-Detailed instructions for the categorization can be found in the Appendix B.1.
-Figure 4: Required Cultural Knowledge.
-Qualitative Examples
-Fig. 3 presents sampled images, questions, and options along with their concept category and question type.
-4 Experiments
-We conduct experiments to evaluate various VLMs on our constructed dataset.
 Model All Food Beverage Game Celeb.
@@
 Agri.
-InstructBLIP-7B (Dai et al., 2024) 50.84 40.85 42.31 38.46 53.19 40.74 50.67 62.16 51.61 60.55 72.22 instructBLIP-13B (Dai et al., 2024) 55.56 45.77 50.00 46.15 59.57 55.56 54.67 64.86 66.13 60.55 64.81 mPLUG-Owl2-7B (Ye et al., 2023) 48.25 42.25 42.31 30.77 63.83 55.56 48.00 54.05 45.16 49.54 66.67 LLaVA-1.6-7B (Liu et al., 2024) 56.32 43.66 48.08 40.38 57.45 51.85 54.67 67.57 59.68 72.48 72.22 LLaVA-1.6-13B (Liu et al., 2024) 57.08 45.07 53.85 36.54 68.09 40.74 53.33 70.27 66.13 69.72 70.37 InternLM-XC2-7B (Dong et al., 2024) 59.67 50.70 48.08 40.38 65.96 55.56 58.67 64.86 69.35 69.72 75.93 Idefics2-8B (Laurençon et al., 2024) 63.62 51.41 50.00 50.00 74.47 66.67 69.33 75.68 74.19 73.39 62.96 Claude-3-opus (Anthropic, 2024) 70.02 62.68 73.08 59.62 72.34 77.78 74.67 78.38 75.81 67.89 75.93 GPT-4-Turbo (Achiam et al., 2023) 80.82 73.94 80.77 78.85 85.11 85.19 81.33 86.49 85.48 79.82 87.04 Gemini-1.5-Pro (Reid et al., 2024) 81.58 80.28 78.85 71.15 85.11 77.78 82.67 83.78 83.87 84.40 85.19
+InstructBLIP-7B (Dai et al., 2024) 50.84 40.85 42.31 38.46 53.19 40.74 50.67 62.16 51.61 60.55 72.22 instructBLIP-13B (Dai et al., 2024) 55.56 45.77 50.00 46.15 59.57 55.56 54.67 64.86 66.13 60.55 64.81 mPLUG-Owl2-7B (Ye et al., 2023) 48.25 42.25 42.31 30.77 63.83 55.56 48.00 54.05 45.16 49.54 66.67 LLaVA-1.6-7B (Liu et al., 2024) 56.32 43.66 48.08 40.38 57.45 51.85 54.67 67.57 59.68 72.48 72.22 LLaVA-1.6-13B (Liu et al., 2024) 57.08 45.07 53.85 36.54 68.09 40.74 53.33 70.27 66.13 69.72 70.37 InternLM-XC2-7B (Dong et al., 2024) 59.67 50.70 48.08 40.38 65.96 55.56 58.67 64.86 69.35 69.72 75.93 Molmo-7B-D (Deitke et al., 2024) 61.04 58.45 71.15 44.23 68.09 44.44 61.33 70.27 56.45 64.22 68.52 Idefics2-8B (Laurençon et al., 2024) 63.62 51.41 50.00 50.00 74.47 66.67 69.33 75.68 74.19 73.39 62.96 Llama-3.2-11B (Dubey et al., 2024) 68.04 61.27 65.38 50.00 72.34 70.37 72.00 75.68 72.58 69.72 81.48 Claude-3-opus (Anthropic, 2024) 70.02 62.68 73.08 59.62 72.34 77.78 74.67 78.38 75.81 67.89 75.93 GPT-4-Turbo (Achiam et al., 2023) 80.82 73.94 80.77 78.85 85.11 85.19 81.33 86.49 85.48 79.82 87.04 Gemini-1.5-Pro (Reid et al., 2024) 81.58 80.28 78.85 71.15 85.11 77.78 82.67 83.78 83.87 84.40 85.19 GPT-4o (OpenAI, 2024) 89.50 88.73 82.69 86.54 95.74 85.19 90.67 91.89 91.94 91.74 87.04
 Table 2: VLMs Evaluation Results.
@@
 The highest and the second highest scores in each column are highlighted in bold and underlined.
+3.1.4 Human Verification
+Our preliminary studies showed that while the VLM often produced factually correct samples, some did not fully align with the cultural nuances we sought to capture.
+Rather than discarding only incorrect content, our human verification process prioritizes selecting question-option sets that best reflect intended cultural subtleties.
+Although this leads to setting aside numerous plausible samples, it does not imply unreliability; instead, it refines the dataset to ensure deep cultural resonance.
+Approved samples require minimal revision before inclusion.
+This approach emphasizes nuanced cultural alignment and indicates the need for further work to better synchronize VLM outputs with human cultural intentions.
+Figure 4: Required Cultural Knowledge.
+3.2 K-Viscuit
+Statistics
+Table 1 presents detailed statistics of our benchmark dataset.
+Our dataset comprises 657 total examples (237 Type 1 and 420 Type 2 questions) based on 237 unique images across 10 concept categories.
+The average word counts are 10.11 and 15.46 for Type 1 and Type 2 questions respectively, with an overall average of 13.53 words.
+Each question includes four options, totaling 2,628 options with an average length of 1.74 words.
+Fig.
+3 shows the distribution of concept categories in our dataset.
+We also compare our dataset size with CVQA (Romero et al., 2024), a recently proposed cultural VQA benchmark, in Appendix E.
+Required Knowledge to Solve Questions
+To characterize our dataset, we analyze the types of knowledge required to solve questions.
+Following Tong et al.
+(2024), we used GPT-4 to analyze Type 2 questions.
+We provided all Type 2 QA pairs and summarized the required knowledge for each question.
+As shown in Fig.
+4, the analysis reveals diverse cultural elements.
+Detailed categorization instructions are provided in Appendix B.1.
+Qualitative Examples
+Fig. 2 showcases sample images, questions, and options along with their concept categories and question types.
+4 Experiments
+We conduct experiments to evaluate various VLMs on our constructed dataset.
 4.1 Models
-The following open-source VLMs are used for experiments: InstructBLIP-7B/13B (Dai et al., 2024), LLaVA-v1.6-7B/13B (Liu et al., 2024), mPLUG-Owl2-7B (Ye et al., 2023), InternLM-XComposer2-VL-7B (Dong et al., 2024), and Idefics2-8B Laurençon et al.
-(2024).
-We also use the following proprietary models: Claude-3-opus (Anthropic, 2024), GPT-4-Turbo (Achiam et al., 2023), and Gemini-1.5-Pro (Reid et al., 2024).
+The following open-source VLMs are used for experiments: InstructBLIP-7B/13B (Dai et al., 2024), LLaVA-v1.6-7B/13B (Liu et al., 2024), mPLUG-Owl2-7B (Ye et al., 2023), InternLM-XComposer2-VL-7B (Dong et al., 2024), Molmo-7B-D (Deitke et al., 2024), Idefics2-8B Laurençon et al.
+(2024), and Llama-3.2-11B-Vision-Instruct (Dubey et al., 2024).
+We also use the following proprietary models: Claude-3-opus (Anthropic, 2024), GPT-4-Turbo (Achiam et al., 2023), Gemini-1.5-Pro (Reid et al., 2024), and GPT-4o (OpenAI, 2024).
 All models are evaluated in the conventional multiple-choice setup, where a model is prompted to choose its answer from one of four options.
@@
 We use accuracy as an evaluation metric.
+4.2 Results
+Table 2 demonstrates the evaluation results of different VLMs on our dataset.
+We first observe that proprietary models usually show higher accuracies.
+The GPT-4o and Gemini-1.5-Pro achieve the highest and second-highest scores, respectively.
+Among the open-sourced models, Llama-3.2-11B and Idefics2-8B usually perform better than other models.
+The accuracy of models in different question types is shown in Table 3.
+We observe that most models show higher accuracy in Type 2 questions compared to Type 1 questions.
+Regarding these trends, we suspect visual recognition with diverse cultural contexts poses inherent challenges for VLMs.
+Our dataset enables evaluating and identifying such aspects where models can be improved.
 Model Type 1 Type 2 Total
@@
 InternLM-XC2-7B 56.12 61.67 59.67
+Molmo-7B-D 55.27 64.29 61.04
 Idefics2-8B 63.71 63.57 63.62
+Llama-3.2-11B 69.20 67.38 68.04
 Claude-3-opus 69.62 80.24 70.02
@@
 Gemini-1.5-Pro 83.97 70.24 81.58
+GPT-4o 92.41 87.86 89.50
 Table 3: Results on Different Question Types. Type 1 and Type 2 denote question types in visual recognition and visual reasoning, respectively.
-4.2 Results
-Table 2 demonstrates the evaluation results of different VLMs on our dataset.
-We first observe that proprietary models usually show higher accuracies.
-The Gemini-1.5-Pro and GPT-4-Turbo achieve the highest and second-highest scores, respectively.
-Among the open-sourced models, Idefics2 and InternLM-XComposer2-VL usually perform better than other models.
-The accuracy of models in different question types is shown in Table 3.
-We observe that most models show higher accuracy in Type 2 questions compared to Type 1 questions, except for Gemini-1.5-Pro and Idefics2-8B.
-Regarding these trends, we suspect visual recognition with diverse cultural contexts poses inherent challenges for VLMs.
-Our dataset enables the evaluation and identification of such aspects where models can be improved.
-Figure 5: Qualitative examples of selected VLMs (i.e., LLaVA-1.6-13B, InternLM-XC2-7B, GPT-4-Turbo, and Gemini-1.5-Pro) on sampled questions.
-The answer option is highlighted in the underline.
-The correct and incorrect choice of models are highlighted in green and red.
+Model en ko en + ko
+Claude-3-Opus 70.02 65.30 70.32
+GPT-4-Turbo 80.82 75.34 80.37
+Gemini-1.5-Pro 81.58 80.82 83.41
+GPT-4o 89.50 76.56 79.76
+Table 4: Results with Different Input Languages. ko is machine-translated texts. For en+ko, we provide questions and options in both languages to models.
 4.3 Analyses
@@
 GPT-4-Turbo and Claude-3-opus usually do not take the benefits of using Korean texts.
-Model en ko en + ko
-Claude-3-Opus 70.02 65.30 70.32
-GPT-4-Turbo 80.82 75.34 80.37
-Gemini-1.5-Pro 81.58 80.82 83.41
-Table 4: Results with Different Input Languages. ko is machine-translated texts. For en+ko, we provide questions and options in both languages to models.
+Figure 5: Human Evaluation.
+Accuracy comparison between Koreans and non-Koreans on 50 samples of K-Viscuit.
+The performance of selected models on this subset is also displayed.
+The average scores for Koreans and non-Koreans are 80.2 and 47.0, respectively.
+Figure 6: Qualitative examples of selected VLMs (i.e., LLaVA-1.6-13B, InternLM-XC2-7B, GPT-4-Turbo, and Gemini-1.5-Pro) on sampled questions.
+The answer option is highlighted in the underline.
+The correct and incorrect choice of models are highlighted in green and red.
 Human Evaluation on the Benchmark
-To evaluate how well people from different backgrounds perform on K-Viscuit, we conducted human evaluation.
+We conducted a human evaluation to evaluate how well people from different backgrounds perform on K-Viscuit.
 We selected a subset of our dataset by randomly sampling 25 images, each with one type 1 and one type 2 question, totaling 50 questions.
@@
 As depicted in Fig.
-6, the results showed that participants with better knowledge of Korean culture achieved higher accuracy.
-However, even among Koreans, knowledge gaps exist both between individuals and within a single person’s knowledge across different domains, such as history.
-Proprietary VLMs can match or exceed human performance, indicating that utilizing VLMs for generating diverse, culturally nuanced questions is more effective than relying solely on individual human annotators.
+5, the results showed that participants with better knowledge of Korean culture achieved higher accuracy.
+However, even among Koreans, knowledge gaps exist between individuals and within a single person’s knowledge across different domains, such as history.
+A Proprietary VLM (i.e., GPT-4-Turbo) shows comparable performance to human performance, indicating that utilizing VLMs for generating diverse, culturally nuanced questions is more effective than relying solely on individual human annotators.
 Details of the user study are provided in the Appendix B.2.
-Figure 6: Human Evaluation. Accuracy comparison between Koreans and non-Koreans on a K-Viscuit subset, measured across different proficiency levels.
+VQAv2 CVQA K-Viscuit
+LLaVA-v1.6-13B 82.8 57.9 57.1
+Molmo-7B-D 85.6 65.5 61.0
+Idefics2-8b 81.2 69.0 63.6
+Llama-3.2-11B 75.2 72.4 68.0
+Table 5: Evaluation results of selected open-source VLMs on various VQA datasets.
+The results on the VQAv2 (Goyal et al., 2017) are taken from the respective papers of each model.
+The performance on the CVQA dataset was measured on the Korean subset.
+Comparison of open-source VLMs on Various VQA Datasets
+We conducted experiments to measure the performance of VLMs on various VQA datasets.
+To this end, we compared the performance of four open-source VLMs on commonly used VQA benchmarks: the VQAv2 (Goyal et al., 2017) dev-test split, the Korean subset of CVQA (Romero et al., 2024), and our dataset, with results shown in Table 5.
+The VLMs demonstrated their best performance on VQAv2, while showing relatively lower accuracy on CVQA and our dataset.
+This suggests that the cultural questions we collected are indeed challenging for VLMs to solve.
 Qualitative Results
@@
 In the second case, two proprietary models are wrong while open-source models make correct answers.
-The third problem requires the model to recognize the structure in the image as Cheomseongdae and infer that it serves a similar function to the Griffith Observatory in the United States.
+The third problem requires the model to recognize the structure in the image as Cheomseongdae and infer that it serves a function similar to that of the Griffith Observatory in the United States.
 Both proprietary models make correct answers to this example.
 In the final example, all models successfully identify the correct answer about agriculture.
+5 Discussion
+We discuss several components and future directions to build VLMs grounded on diverse cultures.
+5.1 Evaluation beyond Multiple-choice VQA
+Our dataset is designed as a multiple-choice Visual Question Answering (VQA) task, where Vision-Language Models (VLMs) select one answer from given options.
+While this classification setup enables straightforward performance measurement through accuracy, it may not fully reveal the models’ depth of cultural understanding.
+For instance, VLMs may correctly select an answer from the available options but fail in a generative VQA setup, where they must generate a free-form answer.
+As shown in Fig.
+7, while the VLM accurately identifies the object as bibimbap in both multiple-choice and generative setups for the first example, it fails to provide accurate cultural context in the second example despite choosing the correct option in the multiple-choice format.
+To quantitatively analyze this aspect, we randomly sampled 80 questions from the Food category and evaluated Llava-v1.6-13B’s performance in a generative setting.
+The model was prompted with the instruction: "This is a question about Korean culture.
+Answer the given questions briefly and concisely." without access to the original multiple-choice options.
+Two native Korean evaluators assessed the generated responses using three categories: (1) Correct: The model’s response fully contained the original answer or was reasonably accurate to the question, (2) Hallucinated: The model confidently provided incorrect information, and (3) Abstained: The model’s response did not include a direct answer to the question but provided a relevant description of the image and did not attempt to generate an answer hastily (e.g., "The specific type of food and its cultural context would determine when it is commonly served").
+The human analysis results are shown in Table 6.
+We observed that the VLM’s accuracy in the generative setup was relatively lower than in the multiple-choice setup (from 45.07% to 36.25%).
+This decrease in accuracy is likely because the task is inherently more difficult without predefined options.
+Interestingly, a significant proportion of responses (21.25%) were classified as abstained, indicating that the VLM recognized the need for additional information to answer confidently and refrained from providing a definitive but potentially incorrect response.
+This behavior suggests that providing the necessary external knowledge could enhance performance.
+Our future research aims to propose various evaluation methods to assess whether VLMs genuinely possess the knowledge to answer questions accurately.
 Figure 7: Qualitative results of a VLM with multi-choice and generative VQA setups.
@@
 The correct and incorrect model generations are manually highlighted by the authors.
-5 Discussion
-We discuss several components and future directions to build VLMs grounded on diverse cultures.
-5.1 Evaluation beyond Multiple-choice VQA
-Our dataset is designed as a multiple-choice VQA task, where VLMs are required to choose one of the given options.
-While this classification setup is advantageous for measuring model performance through accuracy, it may not fully capture the models’ understanding of Korean culture.
-For instance, VLMs may correctly select an answer from the given options but fail in a generative VQA setup where they must generate a free-form answer.
-An example of this is shown in Fig.
-7 (left), where the VLM correctly identifies an object from the choices but hallucinates it as “stones or bricks” when generating an answer without options.
-In contrast, the right example in Fig.
-7 shows the VLM accurately identifying the object as “bibimbap” in both multiple-choice and generative setups.
-Our future research aims to propose various evaluation methods to assess whether VLMs genuinely possess the knowledge to answer questions accurately.
+Correct Hallucinated Abstained Total 29 (36.35%) 34 (42.5%) 17 (21.25%) 80 (100%)
+Table 6: Human evaluation results of LLaVA-1.6-13B on Food category.
 5.2 Augmenting VLMs with External Knowledge
-In previous experiments, we observed significant performance gaps in models for questions that require cultural knowledge.
-To address these gaps, two methodologies can be considered: fine-tuning open-source models and leveraging external knowledge.
-While fine-tuning is limited to open-source models, leveraging external knowledge can be applied to both open-source and proprietary models.
-Therefore, we focused on the latter approach.
+In previous experiments, models struggled with questions requiring cultural knowledge.
+To address these gaps, we considered fine-tuning open-source models and augmenting models with external knowledge.
+Since the latter can be applied to both open-source and proprietary models, we focused on that approach.
 We augmented the models with relevant documents for each test image from the Food concept.
 Our retrieval method involved generating captions using GPT-4-Turbo, embedding these captions via text-embedding-3-large to build queries, and retrieving Top-1 document from 152 Wikipedia pages related to the objects mentioned in the Type 1 options.
-The distractors in K-Viscuit were designed to be highly similar to the correct answers.
-This creates a challenging retrieval setting with numerous hard negatives, aiming to mimic real-world conditions.
-We aim to assess whether the retrieval method remains effective even in such a setting by observing the performance changes when cultural knowledge is added.
-The results show that providing retrieved documents can improve model performance, as indicated by higher scores in the Retrieved setting compared to None.
-The Oracle setting, which uses a curated relevant document, serves as an upper bound.
-Overall, augmenting VLMs with external knowledge via document retrieval demonstrates the potential to bridge the performance gap in VQA tasks requiring cultural context.
+The K-Viscuit distractors were designed to closely resemble correct answers, creating a challenging retrieval setting with numerous hard negatives that mimic real-world conditions.
+We assess if the retrieval method remains effective under these conditions by examining performance changes when cultural knowledge is introduced.
+As shown in Table 7, Providing retrieved documents can enhance model performance, as demonstrated by improved scores in the Retrieved setting compared to None.
+In cases where proprietary models did not benefit from certain retrieved documents, their performance still surpassed the baseline when given carefully curated Oracle documents.
+This suggests that retrieval-based augmentation has the potential to strengthen cultural knowledge in VQA tasks, provided that the quality and relevance of the selected documents are refined.
 We provide the details in the Appendix D.
-Model None Retrieved Oracle
-LLaVA-1.6-7B 43.66 68.31 78.87
-LLaVA-1.6-13B 45.77 64.08 80.28
-InternLM-XC2-7B 50.70 68.31 82.39
-Idefics2-8B 51.41 67.61 82.39
-Claude-3-opus 62.68 70.42 87.32
-GPT-4-Turbo 73.94 78.17 88.73
-Gemini-1.5-Pro 80.28 78.17 90.85
+Model None Retrieved Oracle LLaVA-1.6-7B 43.66 68.31 78.87 LLaVA-1.6-13B 45.77 64.08 80.28 InternLM-XC2-7B 50.70 68.31 82.39 Idefics2-8B 51.41 67.61 82.39 Claude-3-opus 62.68 70.42 87.32 GPT-4-Turbo 73.94 78.17 88.73 Gemini-1.5-Pro 80.28 78.17 90.85 GPT-4o 88.73 83.10 92.25
 Table 7: Retrieval-augmented Generation Results.
 6 Conclusion
-In this work, we proposed a semi-automated pipeline to construct culturally aware benchmarks for vision-language models, featuring multi-choice QA with two types of questions that require visual recognition or reasoning.
-This pipeline combines human-VLM collaboration, enhancing efficiency and diversity in generating culturally relevant QAs.
-Our approach, first used to create the K-Viscuit dataset, can be adapted to other cultures.
-Our evaluation with K-Viscuit identified a notable performance gap between open-source and proprietary VLMs in understanding Korean culture, highlighting the need for improvement in open-source models.
-Further analyses demonstrated the impact of in-depth cultural understanding on VQA results.
-Finally, we enhanced VLM performance through knowledge augmentation and explored open-ended answer generation beyond the multiple-choice setting.
-Our work stresses the importance of cultural diversity in evaluation, aiming for more inclusive and accurate VLMs in the future.
+In this work, we proposed a semi-automated framework for constructing culturally aware benchmarks for vision-language models through human-VLM collaboration, focusing on visual recognition and reasoning in multi-choice QA.
+We demonstrated its effectiveness by creating K-Viscuit, revealing a significant performance gap between proprietary and open-source VLMs in understanding Korean culture.
+Through detailed analyses and knowledge augmentation experiments, we established the impact of cultural understanding on VQA performance and extended our investigation to open-ended answer generation.
+This work highlights the importance of cultural diversity in model evaluation and paves the way for more inclusive VLMs.
 Limitations
@@
 Developing a quality estimation module for generated questions could assist in this process by reducing the workload on human annotators.
+Ethical Statement
+In constructing K-Viscuit, we ensured all images were sourced from Wikimedia Commons under Creative Commons licenses, maintaining proper attribution and copyright compliance.
+The dataset was carefully curated to avoid harmful or inappropriate content, and all questions and answers were verified by native Korean speakers to ensure cultural relevance.
+While comprehensive, we acknowledge that our dataset captures only a subset of Korean cultural elements.
 Appendix A Dataset Construction Details
@@
 Table 12: Prompt for generating image captions.
+Appendix E Comparison of K-Viscuit with CVQA dataset
+Recent advances in vision-language models have sparked a growing interest in evaluating their cultural awareness across diverse global contexts.
+A notable contribution in this direction is CVQA (Romero et al., 2024), which introduces a comprehensive multilingual VQA benchmark spanning 28 countries with 9,044 total examples, averaging approximately 323 examples per country.
+Our K-Viscuit dataset, while focused solely on Korean culture, contains 657 examples—substantially exceeding CVQA’s per-country average.
+Moreover, CVQA’s Korean subset comprises 290 examples, less than half the size of K-Viscuit.
+This focused approach enables K-Viscuit to provide more comprehensive coverage of Korean cultural elements, offering deeper insights into VLMs’ understanding of specific cultural contexts.

### v3

@@
 We also present a series of further analyses, including human evaluation, augmenting VLMs with external knowledge, and the evaluation beyond multiple-choice QA.
-[^1] Our dataset is available at https://huggingface.co/datasets/ddehun/k-viscuit.
+Our dataset is available at https://huggingface.co/datasets/ddehun/k-viscuit.
 Evaluating Visual and Cultural Interpretation:
@@
 K-Viscuit Benchmark with Human-VLM Collaboration
-Yujin Baek11footnotemark: 11     ChaeHun Parkthanks:    Equal contribution1    Jaeseok Kim2     Yu-Jung Heo2    Du-Seong Chang2     Jaegul Choo1 1 KAIST AI     2 KT Corporation {yujinbaek,ddehun,jchoo}@kaist.ac.kr {jaeseok.kim,yj.heo,dschang}@kt.com
+ChaeHun Park11footnotemark: 11     Yujin Baekthanks:    Equal contribution1    Jaeseok Kim2     Yu-Jung Heo2    Du-Seong Chang3     Jaegul Choo1 1 KAIST AI     2 KT Corporation     3 Sogang Univ.
+{ddehun,yujinbaek,jchoo}@kaist.ac.kr {jaeseok.kim,yj.heo}@kt.com     duseong.chang@gmail.com
 1 Introduction
 Recent advances in vision-language models (VLMs) have demonstrated remarkable capabilities in tasks ranging from image captioning to visual question answering.
-However, these models are predominantly trained on Western-centric datasets (Lin et al., 2014; Young et al., 2014; Antol et al., 2015), leading to significant performance disparities when applied to non-Western contexts (Liu et al., 2021; Yin et al., 2021, 2023; Romero et al., 2024).
+However, these models are predominantly trained on Western-centric datasets (Lin et al., 2014; Young et al., 2014; Antol et al., 2015), leading to substantial performance disparities when applied to non-Western contexts (Liu et al., 2021; Yin et al., 2021, 2023; Romero et al., 2024).
 This cultural bias is particularly problematic as visual interpretation often depends heavily on cultural context, necessitating the development of more culturally aware VLMs.
-Several benchmarks have been proposed to evaluate cultural understanding in VLMs (Liu et al., 2021; Yin et al., 2021; Romero et al., 2024; Nayak et al., 2024).
+Figure 1: Dataset Examples. We present an image and two questions of different types for each concept category.
+Several benchmarks have been proposed to evaluate cultural understanding in VLMs (Liu et al., 2021; Yin et al., 2021; Romero et al., 2024; Nayak et al., 2024; Bhatia et al., 2024).
 These approaches primarily rely on manual question generation, which, while valuable, faces certain practical challenges.
 The manual process can be time-consuming and resource-intensive when scaling to new cultural contexts.
-Additionally, as noted in cognitive science research (Ramos, 2020), human annotators may experience cognitive fixation, potentially limiting the diversity of generated questions.
+Additionally, as noted in cognitive science research (Ramos, 2020), humans may experience cognitive fixation, potentially limiting the diversity of generated questions.
 These practical considerations motivate the need for more efficient benchmark construction approaches.
-Figure 1: Framework Overview.
-Inspired by recent successes in human-LLM collaborative data generation (Liu et al., 2022; Bartolo et al., 2022; Kamalloo et al., 2023), we propose a semi-automated framework for constructing cultural VLM benchmarks that enhances both the efficiency and diversity of culture-relevant visual question and answer generation, as shown in Fig.
-1.
+Inspired by recent successes in human-LLM collaborative data generation (Liu et al., 2022; Bartolo et al., 2022; Kamalloo et al., 2023), we propose a semi-automated framework for constructing cultural VLM benchmarks that enhances both the efficiency and diversity of culture-relevant visual question and answer generation (see Fig.
+2).
 Our framework incorporates human-VLM collaboration, where the VLM generates and recommends questions and answers based on carefully crafted guidelines, a small set of human-annotated examples, and image-specific knowledge.
 Native speakers then verify these recommended questions to ensure quality and cultural relevance.
-Using this framework, we develop K-Viscuit (Korean Visual and Cultural Interpretation Test), a benchmark dataset for Korean culture that can be adapted for other cultural contexts.
+Using this framework, we develop K-Viscuit (Korean Visual and Cultural Interpretation Test), a benchmark dataset tailored to Korean culture.
+Note that the framework is adaptable and can be similarly applied to other cultures.
 K-Viscuit features two distinct evaluation types: visual recognition and visual reasoning.
 Also, the benchmark employs carefully designed multiple-choice questions with highly similar distractors to prevent models from exploiting superficial patterns.
-Our evaluation with K-Viscuit reveals a significant performance gap between open-source and proprietary VLMs in understanding Korean culture.
+Our evaluation with K-Viscuit reveals a notable performance gap between open-source and proprietary VLMs in understanding Korean culture.
 We provide insights into the current limitations and potential improvements in VLMs’ cultural understanding capabilities through detailed analyses, which include human evaluation, external knowledge integration, and extended evaluation beyond a multi-choice question-answering setup.
@@
   • We present comprehensive experimental results and analyses of both open-source and proprietary VLMs evaluated on K-Viscuit.
+Figure 2: Framework Overview.
+We begin with image selection, where concept-relevant images (e.g., the traditional Korean game Yutnori) are chosen and briefly described.
+In the QA generation stage, a VLM is guided by annotation guidelines and demonstration samples to draft question–option pairs.
+Finally, human verification ensures annotator review, correction, and filtering of the generated samples.
 2 Related Work
@@
 Our work extends this paradigm to multimodal cultural benchmarks, leveraging AI models to enhance dataset diversity while maintaining high quality through human verification.
-Figure 2: Dataset Examples. We present an image and two questions of different types for each concept category.
 3 Data Construction Framework
@@
 Our dataset encompasses ten core concepts: Food, Beverage, Game, Celebrations, Religion, Tool, Clothes, Heritage, Architecture, and Agriculture.
+Further details are in Appendix B.1.
 3.1.2 Image Selection
@@
 Annotators were encouraged to source diverse and suitable images from various web resources.
+[^1] Although framed as image selection, this step also served as a concept-level filtering process.
+For instance, we excluded abstract or symbolic visuals (e.g., fireworks emojis for “celebration”) that lacked concrete visual referents.
 Wikimedia Commons [^2] https://commons.wikimedia.org/wiki served as the primary source, and only CC-licensed images were selected.
 3.1.3 Question Generation
-Question Type
-Based on the selected images, we annotate questions in multiple-choice QA format.
-To comprehensively evaluate understanding of Korean culture, we categorize questions into two types: visual recognition (Type 1) and reasoning (Type 2).
-Visual recognition questions assess basic visual information such as object identification, while reasoning questions require fine-grained cultural knowledge or deeper reasoning processes related to the image.
-For each image, we created one Type 1 question and between one to four Type 2 questions.
-This categorization offers two key advantages: First, Type 1 questions enable assessment of a model’s basic visual understanding of culturally embedded concepts.
-Second, Type 2 questions comprehensively evaluate cultural understanding beyond simple object recognition.
+Question Type Based on the selected images, we annotate questions in a multiple-choice QA format.
+To comprehensively evaluate the understanding of Korean culture, we categorize questions into two types: Visual Recognition (Type 1) and Cultural Knowledge Application (Type 2).
+Type 1 questions assess basic visual information, such as object identification.
+In contrast, Type 2 questions require a deeper understanding of cultural context, reasoning, or multi-step inference based on the image.
+For each image, we created one Type 1 question and between one and four Type 2 questions.
+This categorization provides two key advantages: First, Type 1 questions assess a model’s ability to recognize culturally embedded visual elements.
+Second, Type 2 questions evaluate cultural understanding beyond simple object recognition.
 # of samples 657
@@
 Table 1: Dataset statistics. The length of questions and options denotes the number of words.
-AI-assisted Question Annotation
-We create questions and their options (one correct answer and three distractors) by leveraging a powerful proprietary VLM (GPT-4-Turbo).
+Figure 3: Concept distribution of our dataset.
+AI-assisted Question Annotation We create questions and their options (one correct answer and three distractors) by leveraging a powerful proprietary VLM (GPT-4-Turbo).
 For each concept category, human annotators first create exemplar questions and options for at least three images.
@@
 (2023), our guidelines emphasize maintaining high similarity among all four multiple-choice options, a principle also reflected in human-annotated examples.
-All information is provided to the VLM through natural language prompts, with distinct annotation guidelines for visual recognition and reasoning questions.
-The detailed prompts are presented in Appendix A.
+All information is provided to the VLM through natural language prompts, with distinct annotation guidelines for visual recognition and application questions.
+The detailed prompts are presented in Appendix B.
+Figure 4: Required Cultural Knowledge.
 It should be noted that all text in the dataset is written in English to isolate the evaluation of multicultural comprehension from multilingual aspects.
@@
 In such cases, we romanize these Korean terms following standard transliteration rules.
-Figure 3: Concept distribution of our dataset.
 Model All Food Beverage Game Celeb.
@@
 Agri.
-InstructBLIP-7B (Dai et al., 2024) 50.84 40.85 42.31 38.46 53.19 40.74 50.67 62.16 51.61 60.55 72.22 instructBLIP-13B (Dai et al., 2024) 55.56 45.77 50.00 46.15 59.57 55.56 54.67 64.86 66.13 60.55 64.81 mPLUG-Owl2-7B (Ye et al., 2023) 48.25 42.25 42.31 30.77 63.83 55.56 48.00 54.05 45.16 49.54 66.67 LLaVA-1.6-7B (Liu et al., 2024) 56.32 43.66 48.08 40.38 57.45 51.85 54.67 67.57 59.68 72.48 72.22 LLaVA-1.6-13B (Liu et al., 2024) 57.08 45.07 53.85 36.54 68.09 40.74 53.33 70.27 66.13 69.72 70.37 InternLM-XC2-7B (Dong et al., 2024) 59.67 50.70 48.08 40.38 65.96 55.56 58.67 64.86 69.35 69.72 75.93 Molmo-7B-D (Deitke et al., 2024) 61.04 58.45 71.15 44.23 68.09 44.44 61.33 70.27 56.45 64.22 68.52 Idefics2-8B (Laurençon et al., 2024) 63.62 51.41 50.00 50.00 74.47 66.67 69.33 75.68 74.19 73.39 62.96 Llama-3.2-11B (Dubey et al., 2024) 68.04 61.27 65.38 50.00 72.34 70.37 72.00 75.68 72.58 69.72 81.48 Claude-3-opus (Anthropic, 2024) 70.02 62.68 73.08 59.62 72.34 77.78 74.67 78.38 75.81 67.89 75.93 GPT-4-Turbo (Achiam et al., 2023) 80.82 73.94 80.77 78.85 85.11 85.19 81.33 86.49 85.48 79.82 87.04 Gemini-1.5-Pro (Reid et al., 2024) 81.58 80.28 78.85 71.15 85.11 77.78 82.67 83.78 83.87 84.40 85.19 GPT-4o (OpenAI, 2024) 89.50 88.73 82.69 86.54 95.74 85.19 90.67 91.89 91.94 91.74 87.04
+InstructBLIP-7B (Dai et al., 2024) 50.84 40.85 42.31 38.46 53.19 40.74 50.67 62.16 51.61 60.55 72.22 instructBLIP-13B (Dai et al., 2024) 55.56 45.77 50.00 46.15 59.57 55.56 54.67 64.86 66.13 60.55 64.81 mPLUG-Owl2-7B (Ye et al., 2023) 48.25 42.25 42.31 30.77 63.83 55.56 48.00 54.05 45.16 49.54 66.67 LLaVA-1.6-7B (Liu et al., 2024a) 56.32 43.66 48.08 40.38 57.45 51.85 54.67 67.57 59.68 72.48 72.22 LLaVA-1.6-13B (Liu et al., 2024a) 57.08 45.07 53.85 36.54 68.09 40.74 53.33 70.27 66.13 69.72 70.37 InternLM-XC2-7B (Dong et al., 2024) 59.67 50.70 48.08 40.38 65.96 55.56 58.67 64.86 69.35 69.72 75.93 Molmo-7B-D (Deitke et al., 2024) 61.04 58.45 71.15 44.23 68.09 44.44 61.33 70.27 56.45 64.22 68.52 Idefics2-8B (Laurençon et al., 2024) 63.62 51.41 50.00 50.00 74.47 66.67 69.33 75.68 74.19 73.39 62.96 Llama-3.2-11B (Dubey et al., 2024) 68.04 61.27 65.38 50.00 72.34 70.37 72.00 75.68 72.58 69.72 81.48 Claude-3-opus (Anthropic, 2024) 70.02 62.68 73.08 59.62 72.34 77.78 74.67 78.38 75.81 67.89 75.93 GPT-4-Turbo (Achiam et al., 2023) 80.82 73.94 80.77 78.85 85.11 85.19 81.33 86.49 85.48 79.82 87.04 Gemini-1.5-Pro (Reid et al., 2024) 81.58 80.28 78.85 71.15 85.11 77.78 82.67 83.78 83.87 84.40 85.19 GPT-4o (OpenAI, 2024) 89.50 88.73 82.69 86.54 95.74 85.19 90.67 91.89 91.94 91.74 87.04
 Table 2: VLMs Evaluation Results.
@@
 3.1.4 Human Verification
-Our preliminary studies showed that while the VLM often produced factually correct samples, some did not fully align with the cultural nuances we sought to capture.
+In this step, we evaluate whether the generated QA pairs adhere not only to the detailed prompt guidelines (Tables 10 and 11), but also to the intended cultural context.
+Our preliminary studies revealed that although VLMs often produced factually accurate outputs, some failed to capture the nuanced cultural elements we aimed to represent.
 Rather than discarding only incorrect content, our human verification process prioritizes selecting question-option sets that best reflect intended cultural subtleties.
 Although this leads to setting aside numerous plausible samples, it does not imply unreliability; instead, it refines the dataset to ensure deep cultural resonance.
-Approved samples require minimal revision before inclusion.
-This approach emphasizes nuanced cultural alignment and indicates the need for further work to better synchronize VLM outputs with human cultural intentions.
-Figure 4: Required Cultural Knowledge.
+In particular, for Type 2 questions, we removed those that primarily tested simple visual recognition rather than cultural knowledge application.
+Approved samples require minimal revision before inclusion, emphasizing nuanced cultural alignment and highlighting the need for further work to better synchronize VLM outputs with human cultural intentions.
 3.2 K-Viscuit
-Statistics
-Table 1 presents detailed statistics of our benchmark dataset.
+Statistics Table 1 presents detailed statistics of our benchmark dataset.
 Our dataset comprises 657 total examples (237 Type 1 and 420 Type 2 questions) based on 237 unique images across 10 concept categories.
@@
 3 shows the distribution of concept categories in our dataset.
-We also compare our dataset size with CVQA (Romero et al., 2024), a recently proposed cultural VQA benchmark, in Appendix E.
-Required Knowledge to Solve Questions
-To characterize our dataset, we analyze the types of knowledge required to solve questions.
+We also compare our dataset size with CVQA (Romero et al., 2024), a recently proposed cultural VQA benchmark, in Appendix A.2.
+Required Knowledge to Solve Questions To characterize our dataset, we analyze the types of knowledge required to solve questions.
 Following Tong et al.
@@
 4, the analysis reveals diverse cultural elements.
-Detailed categorization instructions are provided in Appendix B.1.
-Qualitative Examples
-Fig. 2 showcases sample images, questions, and options along with their concept categories and question types.
+Detailed categorization instructions are provided in Appendix C.1.
+Qualitative Examples Fig. 1 showcases sample images, questions, and options along with their concept categories and question types.
 4 Experiments
-We conduct experiments to evaluate various VLMs on our constructed dataset.
+This section presents our experiments, including VLM evaluation, human benchmarking, and the impact of question language on accuracy.
 4.1 Models
-The following open-source VLMs are used for experiments: InstructBLIP-7B/13B (Dai et al., 2024), LLaVA-v1.6-7B/13B (Liu et al., 2024), mPLUG-Owl2-7B (Ye et al., 2023), InternLM-XComposer2-VL-7B (Dong et al., 2024), Molmo-7B-D (Deitke et al., 2024), Idefics2-8B Laurençon et al.
+The following open-source VLMs are used for experiments: InstructBLIP-7B/13B (Dai et al., 2024), LLaVA-v1.6-7B/13B (Liu et al., 2024a), mPLUG-Owl2-7B (Ye et al., 2023), InternLM-XComposer2-VL-7B (Dong et al., 2024), Molmo-7B-D (Deitke et al., 2024), Idefics2-8B Laurençon et al.
 (2024), and Llama-3.2-11B-Vision-Instruct (Dubey et al., 2024).
@@
 We use accuracy as an evaluation metric.
-4.2 Results
+4.2 Results and Analyses
+Main Results
 Table 2 demonstrates the evaluation results of different VLMs on our dataset.
@@
 Regarding these trends, we suspect visual recognition with diverse cultural contexts poses inherent challenges for VLMs.
+[^3] Further analysis of this trend is provided in A.3.
 Our dataset enables evaluating and identifying such aspects where models can be improved.
-Model Type 1 Type 2 Total
+Model Type 1 Type 2 All
 InstructBLIP-7B 45.57 53.81 50.84
@@
 GPT-4o 92.41 87.86 89.50
-Table 3: Results on Different Question Types. Type 1 and Type 2 denote question types in visual recognition and visual reasoning, respectively.
+Table 3: Results on Different Question Types. Type 1 and Type 2 denote question types in visual recognition and cultural knowledge application, respectively.
+Figure 5: Human Evaluation.
+Accuracy comparison between Koreans and non-Koreans on 50 samples of K-Viscuit.
+The performance of selected models on this subset is also displayed.
+The average scores were 80.2 (Korean participants) and 47.0 (non-Korean participants), with standard deviations of 2.69 and 5.95.
+Human Evaluation on the Benchmark
+We conducted a human evaluation to evaluate how well people from different backgrounds perform on K-Viscuit.
+We selected a subset of our dataset by randomly sampling 25 images, each with one type 1 and one type 2 question, totaling 50 questions.
+This test was administered to 20 Koreans and 14 non-Koreans.
+As depicted in Fig.
+5, the results showed that participants with better knowledge of Korean culture achieved higher accuracy.
+However, even among Koreans, knowledge gaps exist between individuals and within a single person’s knowledge across different domains, such as history.
+A Proprietary VLM (i.e., GPT-4-Turbo) shows comparable performance to human performance, indicating that utilizing VLMs for generating diverse, culturally nuanced questions is more effective than relying solely on individual human annotators.
+Details of the user study are provided in the Appendix C.2.
 Model en ko en + ko
@@
 Table 4: Results with Different Input Languages. ko is machine-translated texts. For en+ko, we provide questions and options in both languages to models.
-4.3 Analyses
+Figure 6: Qualitative examples of selected VLMs (i.e., LLaVA-1.6-13B, InternLM-XC2-7B, GPT-4-Turbo, and Gemini-1.5-Pro) on sampled questions.
+The answer option is highlighted in the underline.
+The correct and incorrect choice of models are highlighted in green and red.
 Asking the VLM in Korean
@@
 GPT-4-Turbo and Claude-3-opus usually do not take the benefits of using Korean texts.
-Figure 5: Human Evaluation.
-Accuracy comparison between Koreans and non-Koreans on 50 samples of K-Viscuit.
-The performance of selected models on this subset is also displayed.
-The average scores for Koreans and non-Koreans are 80.2 and 47.0, respectively.
-Figure 6: Qualitative examples of selected VLMs (i.e., LLaVA-1.6-13B, InternLM-XC2-7B, GPT-4-Turbo, and Gemini-1.5-Pro) on sampled questions.
-The answer option is highlighted in the underline.
-The correct and incorrect choice of models are highlighted in green and red.
-Human Evaluation on the Benchmark
-We conducted a human evaluation to evaluate how well people from different backgrounds perform on K-Viscuit.
-We selected a subset of our dataset by randomly sampling 25 images, each with one type 1 and one type 2 question, totaling 50 questions.
-This test was administered to 20 Koreans and 14 non-Koreans.
-As depicted in Fig.
-5, the results showed that participants with better knowledge of Korean culture achieved higher accuracy.
-However, even among Koreans, knowledge gaps exist between individuals and within a single person’s knowledge across different domains, such as history.
-A Proprietary VLM (i.e., GPT-4-Turbo) shows comparable performance to human performance, indicating that utilizing VLMs for generating diverse, culturally nuanced questions is more effective than relying solely on individual human annotators.
-Details of the user study are provided in the Appendix B.2.
 VQAv2 CVQA K-Viscuit
-LLaVA-v1.6-13B 82.8 57.9 57.1
-Molmo-7B-D 85.6 65.5 61.0
-Idefics2-8b 81.2 69.0 63.6
-Llama-3.2-11B 75.2 72.4 68.0
+LLaVA-v1.6-13B 82.8 57.93 57.08
+Molmo-7B-D 85.6 65.52 61.04
+Idefics2-8b 81.2 68.97 63.62
+Llama-3.2-11B 75.2 72.41 68.04
 Table 5: Evaluation results of selected open-source VLMs on various VQA datasets.
@@
 This suggests that the cultural questions we collected are indeed challenging for VLMs to solve.
+Figure 7: Visual Dependency Analysis.
+Accuracy comparison of selected models when using actual images versus noise images.
+The numbers indicate absolute and relative accuracy drops when replacing images with noise.
+Evaluating the Visual Dependency of K-Viscuit Questions
+To assess how much the K-Viscuit dataset relies on visual information, we conducted an experiment measuring the impact of removing meaningful image content.
+Specifically, models are provided with randomly generated images with pixel values sampled from a Gaussian distribution instead of actual images.
+These images are of the same size as the originals but contain no meaningful visual structure.
+By comparing model performance in this setting to standard image-based evaluation, we can determine the extent to which K-Viscuit questions require visual understanding rather than relying solely on textual cues.
+Fig.
+7 presents the accuracy of selected models under two conditions: with actual images (Image + Text) and with noise images (Text Only).
+Across all models, a substantial drop in accuracy is observed when images are replaced with noise, confirming that K-Viscuit questions require meaningful visual information.
+The performance degradation varies by model, with Llama-3.2-11B showing the largest accuracy drop, while Molmo-7B-D is comparatively less affected.
+Additionally, both Type 1 and Type 2 questions exhibit large performance declines in the absence of images, suggesting that visual information is crucial across different question types.
+These results highlight the varying degrees of visual dependence across models and reinforce the necessity of strong visual reasoning capabilities for solving K-Viscuit questions.
 Qualitative Results
@@
 In the final example, all models successfully identify the correct answer about agriculture.
-5 Discussion
-We discuss several components and future directions to build VLMs grounded on diverse cultures.
+5 Future Directions
+This section explores future directions for improving model evaluation (§5.1) and enhancing model performance with external knowledge (§5.2).
 5.1 Evaluation beyond Multiple-choice VQA
@@
 As shown in Fig.
-7, while the VLM accurately identifies the object as bibimbap in both multiple-choice and generative setups for the first example, it fails to provide accurate cultural context in the second example despite choosing the correct option in the multiple-choice format.
-To quantitatively analyze this aspect, we randomly sampled 80 questions from the Food category and evaluated Llava-v1.6-13B’s performance in a generative setting.
-The model was prompted with the instruction: "This is a question about Korean culture.
-Answer the given questions briefly and concisely." without access to the original multiple-choice options.
-Two native Korean evaluators assessed the generated responses using three categories: (1) Correct: The model’s response fully contained the original answer or was reasonably accurate to the question, (2) Hallucinated: The model confidently provided incorrect information, and (3) Abstained: The model’s response did not include a direct answer to the question but provided a relevant description of the image and did not attempt to generate an answer hastily (e.g., "The specific type of food and its cultural context would determine when it is commonly served").
-The human analysis results are shown in Table 6.
-We observed that the VLM’s accuracy in the generative setup was relatively lower than in the multiple-choice setup (from 45.07% to 36.25%).
-This decrease in accuracy is likely because the task is inherently more difficult without predefined options.
-Interestingly, a significant proportion of responses (21.25%) were classified as abstained, indicating that the VLM recognized the need for additional information to answer confidently and refrained from providing a definitive but potentially incorrect response.
-This behavior suggests that providing the necessary external knowledge could enhance performance.
-Our future research aims to propose various evaluation methods to assess whether VLMs genuinely possess the knowledge to answer questions accurately.
+8, while the VLM accurately identifies the object as bibimbap in both multiple-choice and generative setups for the first example, it fails to provide accurate cultural context in the second example despite choosing the correct option in the multiple-choice format.
 Figure 8: Qualitative results of a VLM with multi-choice and generative VQA setups.
@@
 Correct Hallucinated Abstained Total 29 (36.35%) 34 (42.5%) 17 (21.25%) 80 (100%)
-Table 6: Human evaluation results of LLaVA-1.6-13B on Food category.
+Table 6: Generative QA Results. We manually evaluated the responses of LLaVA-1.6-13B on Food category.
+To quantitatively analyze this aspect, we randomly sampled 80 questions from the Food category and evaluated Llava-v1.6-13B’s performance in a generative setting.
+The model was prompted with the instruction: "This is a question about Korean culture.
+Answer the given questions briefly and concisely." without access to the original multiple-choice options.
+Two native Korean evaluators assessed the generated responses using three categories: (1) Correct: The model’s response fully contained the original answer or was accurate to the question, (2) Hallucinated: The model provided incorrect information, and (3) Abstained: The model’s response did not include a direct answer to the question but provided a relevant description of the image and did not attempt to generate an answer hastily (e.g., "The specific type of food and its cultural context would determine when it is commonly served").
+The human analysis results are shown in Table 6.
+We observed that the VLM’s accuracy in the generative setup was relatively lower than in the multiple-choice setup (from 45.07% to 36.25%).
+This decrease in accuracy is likely because the task is inherently more difficult without predefined options.
+Interestingly, a large proportion of responses (21.25%) were classified as abstained, indicating that the VLM recognized the need for additional information to answer confidently and refrained from providing a definitive but potentially incorrect response.
+This behavior suggests that providing the necessary external knowledge could enhance performance.
+Our future research aims to propose various evaluation methods to assess whether VLMs genuinely possess the knowledge to answer questions accurately.
 5.2 Augmenting VLMs with External Knowledge
@@
 We assess if the retrieval method remains effective under these conditions by examining performance changes when cultural knowledge is introduced.
-As shown in Table 7, Providing retrieved documents can enhance model performance, as demonstrated by improved scores in the Retrieved setting compared to None.
+As shown in Table 7, providing retrieved documents can enhance model performance in the food category, as demonstrated by improved scores in the Retrieved setting compared to None.
 In cases where proprietary models did not benefit from certain retrieved documents, their performance still surpassed the baseline when given carefully curated Oracle documents.
 This suggests that retrieval-based augmentation has the potential to strengthen cultural knowledge in VQA tasks, provided that the quality and relevance of the selected documents are refined.
-We provide the details in the Appendix D.
+We provide the details in the Appendix E.
 Model None Retrieved Oracle LLaVA-1.6-7B 43.66 68.31 78.87 LLaVA-1.6-13B 45.77 64.08 80.28 InternLM-XC2-7B 50.70 68.31 82.39 Idefics2-8B 51.41 67.61 82.39 Claude-3-opus 62.68 70.42 87.32 GPT-4-Turbo 73.94 78.17 88.73 Gemini-1.5-Pro 80.28 78.17 90.85 GPT-4o 88.73 83.10 92.25
-Table 7: Retrieval-augmented Generation Results.
+Table 7: Retrieval-augmented generation results on Food category. Oracle documents were manually selected by the authors.
 6 Conclusion
-In this work, we proposed a semi-automated framework for constructing culturally aware benchmarks for vision-language models through human-VLM collaboration, focusing on visual recognition and reasoning in multi-choice QA.
-We demonstrated its effectiveness by creating K-Viscuit, revealing a significant performance gap between proprietary and open-source VLMs in understanding Korean culture.
-Through detailed analyses and knowledge augmentation experiments, we established the impact of cultural understanding on VQA performance and extended our investigation to open-ended answer generation.
+In this work, we proposed a semi-automated framework for constructing culturally aware benchmarks for vision-language models through human-VLM collaboration, focusing on visual recognition and cultural knowledge application in multi-choice QA.
+We demonstrated its effectiveness by creating K-Viscuit, revealing a performance gap between proprietary and open-source VLMs in identifying Korean cultural concepts.
+Through detailed analyses and knowledge augmentation experiments, we specified the impact of cultural understanding on VQA performance and extended our investigation to open-ended answer generation, discussing its limitations.
 This work highlights the importance of cultural diversity in model evaluation and paves the way for more inclusive VLMs.
@@
 Developing a quality estimation module for generated questions could assist in this process by reducing the workload on human annotators.
+Additionally, LLMs are highly sensitive to the order of answer options in multiple-choice QA tasks, as discussed in Zheng et al.
+(2023); Pezeshkpour and Hruschka (2024).
+While we mitigate positional bias in model evaluation by shuffling the answer options, this approach may not be entirely sufficient.
+As shown in A.1, VLMs often fail to produce consistent predictions when the positions of the answer options are altered.
+Future work could explore calibration techniques, such as contrastive decoding or uncertainty-aware selection strategies, to further enhance robustness in multiple-choice settings.
 Ethical Statement
@@
 While comprehensive, we acknowledge that our dataset captures only a subset of Korean cultural elements.
-Appendix A Dataset Construction Details
-Guidelines to GPT-4-Turbo for Dataset Annotation
+Acknowledgements
+This work was supported by a grant of the KAIST-KT joint research project through AI2X Laboratory, Tech Innovation Group , funded by KT (No.
+D23000019, Research on Multilingual Multicultural Vision-Language Representation Learning) and Institute for Information & communications Technology Planning & Evaluation(IITP) grant funded by the Korea government(MSIT) (RS-2019-II190075, Artificial Intelligence Graduate School Program(KAIST) and RS-2025-02304967, AI Star Fellowship(KAIST).
+Appendix A Additional Analysis
+A.1 Evaluating Sensitivity to Option Order in MCQA with Circular Evaluation
+Recent studies have shown that many vision-language models (VLMs) exhibit sensitivity to the order of answer choices in multiple-choice QA, where performance varies depending on how options are presented.
+To investigate this phenomenon, we apply Circular Evaluation, proposed by Liu et al.
+(2024b).
+This method systematically shifts the positions of four answer choices and forwards them through the model four times, considering a prediction correct only if all variations yield the same answer.
+Figure 9: Order-Sensitivity Analysis.
+Accuracy comparison of selected models under vanilla and circular evaluations (Liu et al., 2024b).
+The numbers indicate the absolute and relative performance drops.
+Fig 9 presents the accuracy of selected models under both vanilla and circular evaluations.
+A significant drop in performance is observed across all models when applying circular evaluation, confirming that answer order affects predictions.
+The relative accuracy decrease ranges from 16.9% to 30.4%, suggesting that models rely on positional cues rather than robust reasoning.
+Among the evaluated models, Molmo-7B-D exhibits the largest performance degradation, indicating heightened sensitivity to ordering effects.
+One potential explanation for this behavior is positional bias, where models develop a preference for certain answer positions.
+These results highlight sensitivity to answer order as a critical challenge in multi-choice QA for vision-language models.
+Future work should explore mitigation strategies, such as answer re-ranking or training objectives that explicitly reduce order sensitivity.
+A.2 Comparison of K-Viscuit with CVQA dataset
+Recent advances in vision-language models have sparked a growing interest in evaluating their cultural awareness across diverse global contexts.
+A notable contribution in this direction is CVQA (Romero et al., 2024), which introduces a comprehensive multilingual VQA benchmark spanning 28 countries with 9,044 total examples, averaging approximately 323 examples per country.
+Our K-Viscuit dataset, while focused solely on Korean culture, contains 657 examples—substantially exceeding CVQA’s per-country average.
+Moreover, CVQA’s Korean subset comprises 290 examples, less than half the size of K-Viscuit.
+This focused approach enables K-Viscuit to provide more comprehensive coverage of Korean cultural elements, offering deeper insights into VLMs’ understanding of specific cultural contexts.
+Upon reviewing the dataset split, we found that 14 answers (4.8%) were simply "Korea," which may limit the flexibility of different prompt settings for evaluation (e.g., "This is an image taken in Korea"), as "Korea" becomes a shortcut response.
+Additionally, CVQA occasionally includes distractors that are inconsistent, such as listing (a) "Gwangju," (b) "Seoul," (c) "Seattle," and (d) "Busan," where (c) "Seattle" is an implausible option.
+This underscores the challenge of ensuring consistency when manually constructing questions and answer choices.
+While CVQA is a valuable concurrent effort, our work focuses on a systematic approach that leverages VLM collaboration to efficiently build culturally specific datasets.
+Category w/o Shot w/ Shot
+Food 72.84 77.78
+Beverage 76.32 76.32
+Game 54.05 56.76
+Celebrations 70.83 83.33
+Religion 70.59 70.59
+Tool 75.56 84.44
+Clothes 81.82 86.36
+Heritage 69.44 91.67
+Architecture 82.50 87.50
+Agriculture 67.50 85.00
+Total 74.31 81.30
+Table 8: Type 2 accuracy with and without one-shot prompting using corresponding Type 1 QA pairs. gpt-4o-mini is used for the experiments.
+A.3 Impact of Type 1 Grounding on Type 2 Reasoning
+In Table 2, we observe that models tend to perform better on Type 2 questions than on Type 1.
+This may seem counterintuitive, as Type 1 questions are grounded directly in visual cues.
+To examine whether grounding in Type 1 content can support Type 2 reasoning, we conducted an experiment where each Type 2 prompt was preceded by its corresponding Type 1 QA pair as a one-shot example.
+As shown in Table 8, this setup led to consistent gains across categories, improving overall accuracy from 74.31% to 81.30%.
+These results suggest visual grounding can meaningfully support higher-level cultural reasoning when leveraged appropriately.
+Core Concept IDS Chapter No.
+Food Food and Beverages (5)
+Beverage Food and Beverages (5)
+Game Motion (10)
+Celebrations Time (14), Religion and belief (22)
+Religion Religion and Belief (19)
+Tool Basic Actions and Technology (9)
+Clothes Clothing and Grooming (6)
+Heritage Cognition (17)
+Architecture The House (7)
+Agriculture Agriculture and Vegetation (8)
+Table 9: Mapping between core concepts and IDS chapters. Some concepts share the same chapter due to overlapping semantic coverage.
+Appendix B Dataset Construction Details
+B.1 Mapping Core Concepts to IDS Chapters
+As described in Section 3.1, following Liu et al.
+(2021), we define the concepts of our dataset based on the Intercontinental Dictionary Series (IDS) (Key and Comrie, 2015).
+Table 9 shows how our 10 core concepts map to chapters in the Intercontinental Dictionary Series (IDS).
+While some mappings are direct, others reflect one-to-many relationships based on semantic grouping and visual realizability (e.g., Food and Beverage both derive from Chapter 5).
+This mapping served as the basis for selecting culturally salient and visually grounded concepts in our dataset.
+B.2 Guidelines to GPT-4-Turbo for Dataset Annotation
 The detailed prompts for the annotation with GPT-4-Turbo are presented in Table 10 and Table 11.
@@
 Answer: Cable Car Distractors: Bus ; Bicycle ; Funicular Railway Please make four options (single answer and three distractors).
-Table 9: A prompt of GPT-4-Turbo used in the annotation of Type 2 questions (i.e., visual reasoning) in the Architecture category.
-Appendix B Dataset Analyses Details
-B.1 Required Knowledge Analysis
+Table 11: A prompt of GPT-4-Turbo used in the annotation of Type 2 questions (i.e., cultural knowledge application) in the Architecture category.
+Appendix C Dataset Analyses Details
+C.1 Required Knowledge Analysis
 We analyzed how diverse the cultural knowledge required by the questions in our K-Viscuit dataset is.
@@
 Table 12: Prompt for required knowledge analysis.
-B.2 Human Evaluation Details
+C.2 Human Evaluation Details
 We randomly selected images according to the proportion of each category to create the questionnaire for the human evaluation.
@@
 We received feedback that Koreans had the most difficulty with questions related to history.
-Appendix C VLM Evaluation Details
+Appendix D VLM Evaluation Details
 Model Implementation Details
@@
 Table 13: Inference prompt for proprietary VLMs.
-Appendix D Retrieval Methodology Details
+Appendix E Retrieval Methodology Details
 For external knowledge retrieval, we generated image captions using GPT-4 with the prompt shown in Table 14.
@@
 Table 14: Prompt for generating image captions.
-Appendix E Comparison of K-Viscuit with CVQA dataset
-Recent advances in vision-language models have sparked a growing interest in evaluating their cultural awareness across diverse global contexts.
-A notable contribution in this direction is CVQA (Romero et al., 2024), which introduces a comprehensive multilingual VQA benchmark spanning 28 countries with 9,044 total examples, averaging approximately 323 examples per country.
-Our K-Viscuit dataset, while focused solely on Korean culture, contains 657 examples—substantially exceeding CVQA’s per-country average.
-Moreover, CVQA’s Korean subset comprises 290 examples, less than half the size of K-Viscuit.
-This focused approach enables K-Viscuit to provide more comprehensive coverage of Korean cultural elements, offering deeper insights into VLMs’ understanding of specific cultural contexts.